### v3

# Super approximation for $\text{SL}_{2}\times\text{SL}_{2}$ and $\text{ASL}_{2}$

Abstract.
Let $S\subset\text{SL}_{2}(\mathbb{Z})\times\text{SL}_{2}(\mathbb{Z})$ or $\text{SL}_{2}(\mathbb{Z})\ltimes\mathbb{Z}^{2}$ be finite symmetric and assume $S$ generates a group $G$ which is a Zariski-dense subgroup $\text{SL}_{2}(\mathbb{Z})\times\text{SL}_{2}(\mathbb{Z})$ or $\text{SL}_{2}(\mathbb{Z})\ltimes\mathbb{Z}^{2}$. We prove that the Cayley graphs
$\{\mathcal{C}ay(G(\operatorname{mod}q),S(\operatorname{mod}q))\}_{q\in\mathbb{Z}}$
form a family of expanders.
Tang and Zhang are supported by ECS grant 27307320, GRF grant 17317222 and NSFC grant 12001457 from the second author.
1. Introduction
Let $G=\langle S\rangle$ be a subgroup of $\text{SL}_{n}(\mathbb{Z})$ with a finite symmetric generating set $S$. For a positive integer $q$, let $G_{q}=G(\operatorname{mod}q)$ and $\Delta_{q}:L^{2}(G_{q})\rightarrow L^{2}(G_{q})$ be defined as, for any $f\in L^{2}(G_{q}),$
$\Delta_{q}(f)(v):=\left(\frac{1}{|S|}\sum_{s\in S}f(\pi_{q}(s)\cdot v)\right)-f(v).$
Each $\Delta_{q}$ is a self-adjoint operator, with spectrum
$0=\lambda_{0,q}>\lambda_{1,q}\geq\lambda_{2,q}\cdots$
For a set $\mathcal{A}$ of positive integers, we say $G$ has super approximation with respect to $\mathcal{A}$ if there is $\epsilon>0$ such that $\lambda_{0,q}-\lambda_{1,q}>\epsilon$, $\forall q\in\mathcal{A}$. The existence of a uniform positive spectral gap only depends on $G$ and is independent of the choice of the finite generating set $S$. If $\mathcal{A}=\mathbb{Z}_{+}$, we simply say $G$ has super approximation.
It is well known that a large class of lattices in $\text{GL}_{n}(\mathbb{Z})$ satisfy the super approximation property, a discovery originally due to Margulis [Mar73] and enhanced later by [BS91] and [Clo03], but the involved techniques (spectral method, Property $T$, automorphic forms, etc.) only works for arithmetic lattices and can not deal with a general non-lattice discrete group. A breakthrough came from Bourgain and Gamburd, who developed the so-called “Bourgain-Gamburd expansion machine”, which is an analytic-combinatorial tool that allows them to prove super approximation property for any Zariski-dense subgroup of $\text{SL}_{2}(\mathbb{Z})$ with respect to prime moduli [BG08b]. A critical ingredient in the proof is Helfgott’s triple product theorem [Hel08]. Since then, there has been a series of papers extending Bourgain-Gamburd’s Theorem to more general groups with respect to more general moduli [BG08a], [BG09], [BGS10], [GV12], [Var12], [BV12], [Gol19], [HdS21].
Regarding the most general possible linear groups having super approximation, in [GV12] Salehi-Golsefidy and Varjú conjectures:
Conjecture 1.1 (Question 2, [GV12]).
Let $G<{\text{SL}}_{n}(\mathbb{Z})$ be finitely generated. Then $G$ has the super approximation property if and only if the identity component $\mathbb{G}_{0}$ of the Zariski closure $\mathbb{G}$ of $G$ is perfect, i.e. $[\mathbb{G}_{0},\mathbb{G}_{0}]=\mathbb{G}_{0}$.
In the same paper [GV12], Salehi-Golsefidy-Varjú proved that $G$ has super approximation with respect to square free numbers if $\mathbb{G}_{0}$ is perfect. Later Salehi-Golsefidy generalized to bounded powers of square free numbers:
Theorem A.
(Salehi-Golsefidy)[Gol19] Let $G<{\text{SL}}_{n}(\mathbb{Z})$ be finitely generated, then $G$ has the super approximation property with respect to bounded powers of square free integers coprime to $q_{0}$ if and only if the identity component $\mathbb{G}_{0}$ of the Zariski closure $\mathbb{G}$ of $G$ is perfect, i.e. $[\mathbb{G}_{0},\mathbb{G}_{0}]=\mathbb{G}_{0}$.
In the paper [Gol19], the author also claims super approximation for $G$ over $\{p^{n}\}_{n\in\mathbb{Z}_{+}}$ for a fixed prime $p$. There appears an issue unless $\mathbb{G}_{0}$ has simple closure. To address this issue, we introduce a gluing technique, which turns out to be applicable to various other settings. We view this gluing technique may be the most novel part of this paper. More detailed discussion is given in Section 2.
If we require no restriction on moduli, all known results have restrictions on the Zariski closure. In [BV12] Bourgain-Varjú proved Conjecture 1.1 when the Zariski closure of $\mathbb{G}_{0}$ is $\text{SL}_{d}$, using an extra ingredient from homogeneous dynamics [BFLM11]. Later de Saxcé and He extended this dynamical tool [HdS19], which allowed them to prove Conjecture 1.1 for all $G$ with simple closure [HdS21].
The full generality of Conjecture 1.1, despite abundant evidence, has still remained a technical challenge. In the paper [Gol20], for the purpose of proving Conjecture 1.1 in general, Salehi-Golsefidy conjectured a sum-product phenomenon over finite quotients of rings of algebraic integers. In a separate paper [TZ23], we proved this conjecture:
Theorem B (Tang-Zhang, 2023).
Suppose $0<\delta\ll 1$, $d$ is a positive integer, and $N_{0}\gg_{d,\delta}1$ is a positive integer. Then there are $0<\varepsilon:=\varepsilon(\delta,d)$ and positive integers $C_{1}=C_{1}(\delta,d),C_{2}=C_{2}(\delta,d),C_{3}=C_{3}(\delta,d)$ such that for any number field $K$ of degree at most $d$ the following holds: Let $\mathcal{O}$ be the ring of integers of $K$. Suppose $\mathfrak{a}$ is an ideal of $\mathcal{O}$ such that $N(\mathfrak{a}):=|\mathcal{O}/\mathfrak{a}|\geq N_{0}$, and suppose $A\subseteq\mathcal{O}$ such that
$|\pi_{\mathfrak{a}}(A)|\geq|\pi_{\mathfrak{a}}(\mathcal{O})|^{\delta}.$
Then there are an ideal $\mathfrak{a}^{\prime}$ of $\mathcal{O}$, and $a\in\mathcal{O}$ such that
$\mathfrak{a}^{C_{1}}\subseteq\mathfrak{a}^{\prime},$
$\pi_{\mathfrak{a}^{\prime}}(\mathbb{Z}a)\subset\pi_{\mathfrak{a}^{\prime}}\left(\sum_{C_{3}}A^{C_{2}}-\sum_{C_{3}}A^{C_{2}}\right),$
$|\pi_{\mathfrak{a}^{\prime}}(\mathbb{Z}a)|\geq N(\mathfrak{a})^{\varepsilon}.$
Here, $\pi_{\mathfrak{a}}\text{ (resp. }\pi_{\mathfrak{a}^{\prime}})$ is the reduction map $\mathcal{O}\rightarrow\mathcal{O}/\mathfrak{a}\text{ (resp. }\mathcal{O}\rightarrow\mathcal{O}/\mathfrak{a}^{\prime})$, the set $A^{C_{2}}$ is the $C_{2}$-fold product of the set $A$, and the set $\sum_{C_{3}}A^{C_{2}}$ is the $C_{3}$-fold sum of the set $A^{C_{2}}$.
Roughly speaking, Conjecture 1.1 is a quantitative statement that given a set $A\subset\mathcal{O}$ and an ideal $\mathfrak{a}\subset\mathcal{O}$, if $|\pi_{\mathfrak{a}}(A)|$ is not too small, then modulo an ideal $\mathfrak{a}^{\prime}$ comparable to $\mathfrak{a}$, a sum-product set of $A$ contains a thick arithmetic progression. With this extra ingredient, we can give a new proof of Conjecture 1.1 in the case that $\text{Zcl}(G)=\text{SL}_{2}$, a result due to Bourgain-Varjú, following the spirit of [BG09]. Combining Theorem B with a gluing technique which we think is our main technical contribution, we can prove Conjecture 1.1 for two representative open cases of Conjecture 1.1, which is the main theorem of this paper:
Theorem 1.2.
Let $S$ be a finite symmetric set that generates a group $G$ which is a Zariski dense subgroup of ${\text{SL}}_{2}(\mathbb{Z})\times{\text{SL}}_{2}(\mathbb{Z})$ or ${\text{SL}}_{2}(\mathbb{Z})\ltimes\mathbb{Z}^{2}$, then $G$ has super approximation property with respect to all positive integers.
Theorem 1.2 confirms Conjecture 1.1 for the first semisimple but non-simple clousre case, and the first non-semisimple closure case.
One advantage of our approach using sum-product, compared to the method in [BV12] and [HdS21] is that we can allow denominators for the group. With little modification of the proof, we have the following extension of Theorem 1.2:
Theorem 1.3.
Let $S$ be a finite symmetric set in ${\text{SL}}_{2}(\frac{1}{q_{0}}\mathbb{Z})\times{\text{SL}}_{2}(\frac{1}{q_{0}}\mathbb{Z})$ $({\text{SL}}_{2}(\frac{1}{q_{0}}\mathbb{Z})\ltimes(\frac{1}{q_{0}}\mathbb{Z})^{2}$, respectively)))), and $S$ generates a group $G$ which is a Zariski dense in ${\text{SL}}_{2}\times{\text{SL}}_{2},({\text{ASL}}_{2}$, respectively)))), then $G$ has super approximation with respect to $\{q\in\mathbb{Z}_{+},(q,q_{0})=1\}$.
Acknowledgements
We thank Zeev Rudnick, Nicolas de Saxcé, He Weikun and Tran Chieu-Minh for many helpful corrections/suggestions on previous versions of this paper.
2. Notations and methodology
We introduce the following notations which we use throughout this paper.
The unit of any multiplicatively written group is denoted by 1. Occasionally, if a ring structure is present, we denote the additive unit by 0. For given two subsets $A$ and $B$, we denote their product set by $A\cdot B=\{ab|a\in A,b\in B\}$, and their sum set by $A+B=\{a+b|a\in A,b\in B\}$. The $k$-fold product of $A$ is denoted by $A^{k}$, and the $k$-fold sum of $A$ is denoted by $\sum_{k}A$.
If $f$ and $g$ are two complex valued functions on a discrete group $G$, we denote by $f*g$ their convolution
$f*g(x)=\sum_{y\in G}f(y)g(xy^{-1}).$
We write $f^{(l)}$ for the $k$-fold convolution of $f$ with itself.
For a prime $p$, we write $p^{n}||q$ if $p^{n}|q$ but $p^{n+1}\nmid q$, and we define $v_{p}(q)=n$. Similarly, for a matrix $M$, $v_{p}(M)=n$ if $p^{n}|M$ but $p^{n+1}\nmid M$. For a prime ideal $\mathcal{P}$ in a ring of integers $\mathcal{O}$, we can define $v_{\mathcal{P}}$ in an analogous way. For two integers $q_{1}$ and $q_{2}$, we write $q_{1}||q_{2}$ if for every $p^{n}||q_{1}$, we also have $p^{n}||q_{2}$. The exact division ``||"``||"` ` | | " can be extended to ideals in a natural way. For a prime ideal $\mathcal{P}$ and a general ideal $\mathfrak{a}$ of a ring $\mathcal{O}$, we write $\mathcal{P}^{n}||\mathfrak{a}$ if $\mathcal{P}^{n}\supset\mathfrak{a}$ but $\mathcal{P}^{n+1}\not\supset\mathfrak{a}$. Similarly, for two ideals $\mathcal{Q}_{1}\supset\mathcal{Q}_{2}$, we write $\mathcal{Q}_{1}||\mathcal{Q}_{2}$ if for all $\mathcal{P}^{n}||\mathcal{Q}_{1}$, we also have $\mathcal{P}^{n}||\mathcal{Q}_{2}$.
For $q=\prod_{i}p_{i}^{n_{i}}\in\mathbb{Z}_{+}$ and $\alpha$ a real positive number, we let $q^{\{\alpha\}}=\prod_{i}p_{i}^{[n_{i}\alpha]}$, where $[n_{i}\alpha]$ is the integer part of $n_{i}\alpha$.
Let $\pi_{q}:\mathbb{Z}\rightarrow q\mathbb{Z}$ be the residue map, which induces residue maps in various other contexts, and we denote them by $\pi_{q}$ as well.
Let $\Lambda=\text{SL}_{2}(\mathbb{Z})\times\text{SL}_{2}(\mathbb{Z})$ or $\text{SL}_{2}(\mathbb{Z})\ltimes\mathbb{Z}^{2}$, $\Lambda_{q}=\Lambda(\operatorname{mod}q)$ and $\Lambda(q)$ be the kernel of the reduction map $\pi_{q}:\Lambda\rightarrow\Lambda_{q}$. Sometimes we need to reduce the two factors of $\Lambda$ by two different moduli $q_{1},q_{2}$, and we denote this reduction by $\pi_{q_{1},q_{2}}$. Throughout our paper, we fix $\Gamma=\text{SL}_{2}(\mathbb{Z})$.
We use the standard asymptotic notations $O,\Omega,\Theta$ to describe growth of functions. For convenience we also adopt Bourgain’s notations: we write $f(q)<q^{c+}$ to mean $f(q)<q^{c+\epsilon}$ for arbitrarily small $\epsilon$ when $q$ large. Similarly, $f(q)>q^{c-}$ to mean $f(q)>q^{c-\epsilon}$ for arbitrarily small $\epsilon$ when $q$ large.
Let $\chi_{S}$ be the normalized uniform counting measure supported on $S$, i.e., for $A\subset G$, $\chi_{S}(A)=\frac{|A\cap S|}{|S|}$. Let $\pi_{q}[\chi_{S}]$ be the pushforward of $\chi_{S}$ under the residue map $\pi_{q}$. Let $T_{q}$ be the convolution operator by $\pi_{q}[\chi_{S}]$, i.e., For $f\in L^{2}(G_{q})$,
$T_{q}(f)=\pi_{q}[\chi_{S}]*f.$ (2.1)
Then $T_{q}$ is a self adjoint operator on $L^{2}(G_{q})$ with an invariant subspace $l_{0}^{2}(G_{q})$ consisting of functions with average 0. Denote the set of eigenvalues of $T_{q}$ on $l_{0}^{2}(G_{q})$ by $E_{q}$. It follows from Alon and Milman [AM85] that $G$ has super approximation with respect to $\mathbb{Z}_{+}$ if and only if there is some constant $c<1$ independent of $q$ such that $\lambda<c$ for all eigenvalues $\lambda\in E_{q}$. Then following the argument in [BV12] (Proof of Theorem 1, Page 156-158), Theorem 1.2 can be derived from the following proposition:
Proposition 2.2.
Let $S$ be a symmetric set in $\Lambda={\text{SL}}_{2}(\mathbb{Z})\times{\text{SL}}_{2}(\mathbb{Z})$ or ${\text{SL}}_{2}(\mathbb{Z})\ltimes\mathbb{Z}^{2}$, and assumle that it generates a group $G$ which is Zariski-dense in ${\text{SL}_{2}\times\text{SL}_{2}}$ or ${\text{ASL}_{2}}$. Then for any $\epsilon>0$ there are $10<\delta<$ 1, $C>0$ depending on $\epsilon$ such that the following holds. If $q\in\mathbb{Z}_{+}$ sufficiently large, $A\subset\Lambda$ symmetric, and some integer $l$ satisfying
$\chi_{S}^{(l)}(A)>q^{-\delta},\hskip 14.22636ptl>\delta^{-1}\log q\hskip 14.22636pt\text{ and }\hskip 14.22636pt|\pi_{q}(A)|<|\Lambda_{q}|^{1-\epsilon},$ (2.3)
then
$|\pi_{q}(A\cdot A\cdot A)|>|\pi_{q}(A)|^{1+\delta}.$ (2.4)
It is noted that the reduction in [BV12] uses bounds for multiplicities of nontrivial irreducible representations appearing in the regular representation of $\text{SL}_{2}(\mathbb{Z}/q\mathbb{Z})$. Similar bounds exist for $\text{SL}_{2}(\mathbb{Z}/q\mathbb{Z})\times\text{SL}_{2}(\mathbb{Z}/q\mathbb{Z})$ and $\text{SL}_{2}(\mathbb{Z}/q\mathbb{Z})\ltimes(\mathbb{Z}/q\mathbb{Z})^{2}$ as well.
In the rest of the paper we will focus on proving Proposition 2.2. We give a sketch here.
2.1. Sketch of proof for Proposition 2.2
We assume (2.4) does not hold for a sufficiently small $\delta$. We will then show a bounded product of $A$ has size at least $|\Lambda_{q}|^{1-\epsilon/2}$, which will force (2.4) to hold afterall.
We assume $\Lambda=\text{SL}_{2}(\mathbb{Z})\times\text{SL}_{2}(\mathbb{Z})$ and we let $\mathbb{P}_{1},\mathbb{P}_{2}$ be the projections of $\Lambda$ to its first and second components. For simplicity we assume each prime divisor of $q$ has sufficiently large exponent. On the first stage, we follow closely the method in [BG08a] and [BG09] and apply Theorem B to show that there is a not-too-small exact divisor $q^{\prime}$ of $q$ and some constant $C$ such that one simple factor, say $\pi_{q^{\prime}}\circ\mathbb{P}_{1}(A^{C})$ contains a large congruence subgroup of $\Gamma_{q^{\prime}}$ (Proposition 5.1). In this process, compared to [BG09], we apply Theorem A to take a shortcut to deduce some results on non-concentration of random walks at Archimedean place.
However, for the purpose of proving Proposition 2.2, we need to take $q^{\prime}$ a very large divisor of $q$, and this leads to the following question:
Suppose for two exact divisors $q_{1},q_{2}$ of $q$, we have a set $B\subset\Gamma_{q_{1}}\times\Gamma_{q_{2}}$ such that $\mathbb{P}_{1}(B),\mathbb{P}_{2}(B)$ are large in $\Gamma_{q_{1}},\Gamma_{q_{2}}$, can we boundedly generate a large subset of $\Gamma_{q_{1}}\times\Gamma_{q_{2}}$ by $B$?
We need to deal with the case $q_{1},q_{2}$ coprime for the above question when growing $q^{\prime}$ to a very large divisor of $q$ for one simple factor, and we also need to consider the case $q_{1},q_{2}$ non co-prime when passing from one simple factor to two simple factors.
In the paper [Gol19], the author also claims super approximation for a general group with perfect Zariski closure with respect to $\{p^{n}\}_{n\in\mathbb{Z}_{+}}$. We are reluctant to cite this result unless the group has simple closure, as there seems to be an issue in the paper [SG17] where this claim is based on, and this issue already exists in the case $\text{SL}_{2}(\mathbb{Z}/p^{n}\mathbb{Z})\times\text{SL}_{2}(\mathbb{Z}/p^{n}\mathbb{Z})$. By first applying sum-product which creates thick one-parameter groups in one simple factor while losing control of the other factor, and then taking commutators, the author essentially arrives at the situation of the aforementioned question for the case $q_{1}=q_{2}=p^{n}$. It appears the author’s argument is that $\mathbb{P}_{1}(B),\mathbb{P}_{2}(B)$ is large implies $B$ is large, for which simple counter examples can be found. See the sentence “Lemma 71 and Lemma 34 imply Proposition 37” (Pg. 7254, [SG17]) and trace from there.
To address this question, our approach is as follows. For simplicity we assume $\mathbb{P}_{1}(B)=\Gamma_{q_{1}}$ and $\mathbb{P}_{2}(B)=\Gamma_{q_{2}}$. We consider a section map
$\psi:\Gamma_{q_{1}}\rightarrow B$
such that $\mathbb{P}_{1}\circ\psi$ is the identity map. Take a small parameter $10<\theta<$ 1, e.g. $\theta=\frac{1}{1000}$. According to a dichotomy (Proposition 3.1), there are two scenarios:
  1. (1)
    There exists a large exact divisor $q_{3}$ of $q_{2}$, and $x,y\in\Gamma_{q_{1}}$ such that for any $p^{n}||q_{3}$, $\psi(xy)\neq\psi(x)\psi(y)(\operatorname{mod}p^{[n\theta]})$.
  2. (2)
    There exists a large exact divisor ${q_{4}}$ of $q_{2}$, and a large subset $S$ of $\Gamma_{q_{1}}$ such that for any $x,y\in S$, any $p^{n}||q_{4}$, $\psi(xy)=\psi(x)\psi(y)(\operatorname{mod}p^{[n\theta]})$.
In the first scenario, one can take $B^{\prime}$ to be the collection of all conjugates of $\psi(xy)\psi(x)^{-1}\psi(y)^{-1}$ by $B$ so that $\mathbb{P}_{1}(B^{\prime})=1$ and $\pi_{q_{3}}\circ\mathbb{P}_{2}(B^{\prime})$ is large, thus $\pi_{q_{1},q_{3}}(BB^{\prime})$ is large in $\Gamma_{q_{1}}\times\Gamma_{q_{3}}$.
In the second scenario, we consider two cases. If for a large exact divisor $q_{4}^{\prime}$ of ${q_{4}}$, we have $\pi_{q_{3}^{\{\theta/2\}}}\circ\mathbb{P}_{2}\circ\psi=1$, then by taking commutator of $\psi(\Gamma_{q_{1}})$ iteratively one can create a set $B^{\prime\prime}$ such that $\pi_{q_{4}^{\prime}}\circ\mathbb{P}_{2}(B^{\prime\prime})=1$ and $\mathbb{P}_{1}(B^{\prime\prime})$ is large in $\Gamma_{q_{1}}$, so that $\pi_{q_{1},q_{4}^{\prime}}(BB^{\prime\prime})$ is large in $\Gamma_{q_{1}}\times\Gamma_{q_{4}^{\prime}}$.
If for a large exact divisor $q_{4}^{\prime\prime}$ of ${q_{4}}$ such that for any $p^{n}||q_{4}^{\prime\prime}$, we have $\pi_{q_{3}^{\{{\theta}/{2}\}}}\circ\mathbb{P}_{2}\circ\varphi\neq 1$, then we must have $q_{4}^{\prime\prime}||q_{1}$, and we can find a set $B^{\prime\prime\prime}\subset\text{Im}(\psi)$ such that $B^{\prime\prime\prime}$ contains all the preimage of a set $\mathcal{P}\subset\Gamma_{Q_{1}^{2}}\times\Gamma_{Q_{2}^{2}}$ under the following canonical projection:
$\Gamma_{q_{1}}\times\Gamma_{q_{4}^{\prime\prime}}\cong\Gamma_{q_{1}/q_{4}^{\prime\prime}}\times\Gamma_{q_{4}^{\prime\prime}}\times\Gamma_{q_{4}^{\prime\prime}}\rightarrow\Gamma_{q_{4}^{\prime\prime}}\times\Gamma_{q_{4}^{\prime\prime}}\rightarrow\Gamma_{Q_{1}^{2}}\times\Gamma_{Q_{2}^{2}}$
where $Q_{1},Q_{2}|q_{4}^{\prime\prime}$ such that for each $p^{n}||q_{4}^{\prime\prime}$, $v_{p}(Q_{1}),v_{p}(Q_{2})=\Theta(n\theta)$, and $\mathcal{P}=\{(1,1)+n(Q_{1}X_{1},Q_{2}X_{2})\}_{n\in\mathbb{Z}}$, where $X_{1},X_{2}$ are two traceless primitive matrices in $\text{Mat}(2,\mathbb{Z})$. Then one can apply quantitative non-concentration of random walks to find elements from $A$ to conjugates $(X_{1},X_{2})$ to different directions. Multiplying these conjugates together and taking commutators, we can recover a large set in $\Gamma_{q_{1}}\times\Gamma_{q_{4}^{\prime\prime}}$.
This whole scheme is carried out in the proof of Proposition 6.1. One can then apply Proposition 6.1 iteratively to create a large subset of $\Lambda_{q}$, which forces (2.4) to hold afterall.
The idea of considering section maps comes from Bourgain’s gluing scheme for $\mathbb{Z}_{q}$ (See Lemma 6.71 of [Bou08]). If the group is non-semismple, this method can also be applied to connect the semisimple part of the group to the unipotent part. See Section 8.
Remark 2.5.
Since Proposition 5.1 is directly given by the main theorem of [BV12], if only for a proof for Theorem 1.2, one can skip Section 5 and start from Section 6, and reference back to more or less standard lemmas in Sections 3 and 4 when necessary. We include a detailed proof of Proposition 5.1 to deal with groups with denominators (Theorem 1.3) and for potential further generalization.
3. Preliminaries on combinatorics
The first ingredient is a generalization of Corollary 6.9 in [Bou08] which proved the case $G_{1},G_{2}$ abelian and $\text{gcd}(|G_{1}|,|G_{2}|)=1$.
Proposition 3.1.
Let $G_{1},G_{2}$ be two finite groups and let $\psi:G_{1}\rightarrow G_{2}$ some map. Then for $0<\varepsilon<\frac{1}{1600}$ we have either
|$\left|\left\{(x,y)\in G_{1}\times G_{1}\mid\psi(xy)=\psi(x)\psi(y)\right\}\right|<\left(1-\varepsilon\right)\left|G_{1}\right|^{2},$ (3.2)
or there exists a subset $S\subset G_{1}$ with $|S|>(1-\sqrt{\varepsilon})|G_{1}|$ and a group homomorphism $f:G_{1}\rightarrow G_{2}$ such that
$\psi|_{S}=f|_{S}.$ (3.3)
Remark 3.4.
If $\text{gcd}\left(\left|G_{1}\right|,\left|G_{2}\right|\right)=1$, then (3.3) implies
|$\left|\left\{x\in G_{1}\mid\psi(x)\neq 1\right\}\right|<\sqrt{\varepsilon}\left|G_{1}\right|.$
Proposition 3.1 relies on the following Lemma 3.5 and Theorem 3.6.
Lemma 3.5.
Let $A$ be a finite subset of a group $Z$ and $\mathcal{G}\subset A\times A$, $40<\varepsilon<1/$ 4, such that
$|\mathcal{G}|>(1-\varepsilon)|A|^{2}$
Then there exists a subset $A^{\prime}$ of $A$ satisfying
$\left|A^{\prime}\right|>(1-\sqrt{\varepsilon})|A|$
and
$\left|A^{\prime}A^{\prime}\right|<\frac{|A\stackrel{{\scriptstyle\mathcal{G}}}{{\cdot}}A|^{4}}{(1-\sqrt{\varepsilon})(1-2\sqrt{\varepsilon})^{2}|A|^{3}},$
where $A\stackrel{{\scriptstyle\mathcal{G}}}{{\cdot}}A:=\{ab\mid a,b\in A,(a,b)\in\mathcal{G}\}$.
Proof of Lemma.
3.5. The proof was stated in [Bou08] for an abelian group $Z$. The same proof works for a general group as well. ∎
Theorem 3.6 (Noncommutative Freiman-Kneser theorem for small doubling).
Let $0<\varepsilon\leq 1$, and let $S\subset G$ be a finite non-empty subset of a multiplicative group $G$ such that $|A\cdot S|\leq(2-\varepsilon)|S|$ for some finite set $A$ of cardinality $|A|$ at least $|S|$, where $A\cdot S:=\{as:a\in A,s\in S\}$ is the product set of $A$ and $S$. Then there exists a finite subgroup $H$ of $G$ with cardinality $|H|\leq C(\varepsilon)|S|$, such that $S$ is covered by at most $C^{\prime}(\varepsilon)$ right-cosets $H\cdot x$ of $H$ for some $C(\varepsilon),C^{\prime}(\varepsilon)\leq{2}/{\varepsilon}-1$.
Remark 3.7.
Theorem 3.6 is due to Hamidoune. See the article “Hamidoune’s Freiman-Kneser theorem for nonabelian groups” in Terence Tao’s blog which gives a concise proof of Theorem 3.6.
Proof of Proposition 3.1.
[…]
$\mathcal{G}=\left\{(x,y)\in G_{1}\times G_{1}\mid\psi(xy)=\psi(x)\psi(y)\right\}$
satisfies
$|\mathcal{G}|\geq\left(1-\varepsilon\right)\left|G_{1}\right|^{2}.$
Denote
$A=\left\{(x,\psi(x))\mid x\in G_{1}\right\}\subset G_{1}\times G_{2}$
and let $\mathcal{G}^{\prime}=\{((x,\psi(x)),(y,\psi(y))):(x,y)\in\mathcal{G}\}$. Then $\mathcal{G}^{\prime}$ is a graph on $A$ and $|\mathcal{G}^{\prime}|=|\mathcal{G}|\geq(1-\epsilon)|G_{1}|^{2}=(1-\epsilon)|A|^{2}$. Apply Lemma 3.5 with $Z=G_{1}\times G_{2}$. We obtain a subset $A^{\prime}\subset A$ satisfying
$\left|A^{\prime}\right|>(1-\sqrt{\varepsilon})\left|G_{1}\right|$ (3.8)
and
$\left|A^{\prime}A^{\prime}\right|<\frac{\left|G_{1}\right|}{\left(1-\sqrt{\varepsilon}\right)\left(1-2\sqrt{\varepsilon}\right)^{2}}<\left(1+10\sqrt{\varepsilon}\right)\left|A^{\prime}\right|<\frac{5}{4}\left|A^{\prime}\right|.$ (3.9)
Next, apply Theorem 3.6 to $A^{\prime}\subset G_{1}\times G_{2}$. There is a subgroup $H$ of $G_{1}\times G_{2}$ and $\left(x_{1},x_{2}\right)\in G_{1}\times G_{2}$ such that
$A^{\prime}$ $\subset\left(x_{1},x_{2}\right)\cdot H,$ (3.10)
$|H|$ $<\frac{5}{3}\left|A^{\prime}\right|.$ (3.11)
[…]
Let $H_{1}=\mathbb{P}_{1}\left(H\right)$. We have
[…]
implying that $H_{1}=G_{1}$. Then for any $x\in G_{1},$ there exists $f(x)\in G_{2}$ such that $(x,f(x))\in H$. Assume there exists $(y,z_{1}),(y,z_{2})\in H$ with $z_{1}\neq z_{2}$. Then for any $x\in G_{1}$, we have $(x,f(x)z_{1}z_{2}^{-1})\in H$ with $(x,f(x)z_{1}z_{2}^{-1})\neq(x,f(x))$. So $|H|\geq 2|G_{1}|\geq 2|A^{\prime}|$. Contradiction. Hence the choice of $f(x)$ is unique for all $x$. Since $H$ is a subgroup, we get $(1,1)\in H$ so $f(1)=1$. Also we see $(y_{1},f(y_{1}))(y_{2}^{-1},f(y_{2}^{-1}))=(y_{1}y_{2}^{-1},f(y_{1})f(y_{2}^{-1}))=(y_{1}y_{2}^{-1},f(y_{1}y_{2}^{-1}))$, so $f$ is a group homomorphism.
Since
$\left|\mathbb{P}_{1}\left(A^{\prime}\right)\right|=|A^{\prime}|>\left(1-\sqrt{\varepsilon}\right)\left|G_{1}\right|,$
we deduce
$\left|\mathcal{G}\cap\left(\mathbb{P}_{1}\left(A^{\prime}\right)\times\mathbb{P}_{1}\left(A^{\prime}\right)\right)\right|\geq|\mathcal{G}|+|\mathbb{P}_{1}\left(A^{\prime}\right)\times\mathbb{P}_{1}\left(A^{\prime}\right)|-|G_{1}|^{2}>\left(1-2\sqrt{\varepsilon}\right)\left|G_{1}\right|^{2}$
by inclusion-exclusion; and hence
$\left|\mathbb{P}_{1}\left(A^{\prime}\right)\stackrel{{\scriptstyle\mathcal{G}}}{{\cdot}}\mathbb{P}_{1}\left(A^{\prime}\right)\right|\geq\frac{\left|\mathcal{G}\cap\left(\mathbb{P}_{1}\left(A^{\prime}\right)\times\mathbb{P}_{1}\left(A^{\prime}\right)\right)\right|}{|G_{1}|}>(1-2\sqrt{\varepsilon})\left|G_{1}\right|.$
Therefore, $\mathbb{P}_{1}\left(A^{\prime}\right)\cap\left[\mathbb{P}_{1}\left(A^{\prime}\right)\stackrel{{\scriptstyle\mathcal{G}}}{{\cdot}}\mathbb{P}_{1}\left(A^{\prime}\right)\right]\neq\emptyset$, and for $\left(y_{1},y_{2}\right)\in\mathcal{G}\cap\left(\mathbb{P}_{1}\left(A^{\prime}\right)\times\mathbb{P}_{1}\left(A^{\prime}\right)\right)$ with $y_{1}y_{2}\in\mathbb{P}_{1}\left(A^{\prime}\right)$, we get
$x_{2}f(x_{1}^{-1}y_{1}y_{2})=\psi\left(y_{1}y_{2}\right)=\psi\left(y_{1}\right)\psi\left(y_{2}\right)=x_{2}f(x_{1}^{-1}y_{1})x_{2}f(x_{1}^{-1}y_{2}).$
$\Rightarrow f(x_{1})=x_{2}\Rightarrow(x_{1},x_{2})\in H\Rightarrow A^{\prime}\subset H.$
We finish the proof of Proposition 3.1 by taking $S=\mathbb{P}_{1}\left(A^{\prime}\right)$. ∎
We also need the following bounded generation result over $\text{SL}_{2}(\mathbb{Z})\times\text{SL}_{2}(\mathbb{Z})$.
Proposition 3.12.
For any $0<\delta<\frac{1}{25}$, the following holds: Let $A\subset\mathrm{SL}_{2}(\mathbb{Z}/q_{1}\mathbb{Z})\times\mathrm{SL}_{2}(\mathbb{Z}/q_{2}\mathbb{Z})$ be symmetric and $|A|>(q_{1}q_{2})^{3-\delta}$. Then there exists $q_{1}^{\prime}|q_{1},q_{2}^{\prime}|q_{2},q_{1}^{\prime}q_{2}^{\prime}<(q_{1}q_{2})^{40\delta}$ such that
$A^{2880}\supset\mathrm{SL}_{2}(q_{1}^{\prime}\mathbb{Z}/q_{1}\mathbb{Z})\times\mathrm{SL}_{2}(q_{2}^{\prime}\mathbb{Z}/q_{2}\mathbb{Z}).$
In [Hel08] Helfgott proved that given $A\subset\text{SL}_{2}(\mathbb{F}_{p})$ with $|A|>p^{\frac{8}{3}}$, then $A\cdot A\cdot A={\text{SL}}_{2}(\mathbb{F}_{p})$ for a sufficiently large prime $p$. Helfgott’s result can be obtained in an elegant way by a representation theoretical approach due to Gowers [Gow08], but Gower’s approach does not seem to adapt well for a general composite modulus. We follow Helfgott’s approach for the proof of Proposition 3.12.
We start with a few lemmas.
Lemma 3.13.
Let $0<\gamma<\frac{1}{4}$ and let $A,B\in\mathbb{Z}/q\mathbb{Z}$ with $|A|,|B|>q^{1-\gamma}$, then there exists $q^{\prime}|q,q^{\prime}<q^{\frac{12\gamma}{5}}$ such that
$q^{\prime}\mathbb{Z}/q\mathbb{Z}\subset\sum_{24}AB-AB.$
Proof.
This is a special case of Corollary A.7 in [TZ23]. ∎
We need a two dimensional analog of Lemma 3.13.
Lemma 3.14.
Let $0<\delta<\frac{1}{8}$. Let $q_{1},q_{2}\in\mathbb{Z}$ and $A,B$ be subsets of $\mathbb{Z}/q_{1}\mathbb{Z}\times\mathbb{Z}/q_{2}\mathbb{Z}$ such that $|A|,|B|>(q_{1}q_{2})^{1-\delta}$. Then there exists $q_{1}^{\prime}|q_{1},q_{2}^{\prime}|q_{2},q_{1}^{\prime}q_{2}^{\prime}<(q_{1}q_{2})^{10\delta}$, such that
$\sum_{96}AB-AB\supset q_{1}^{\prime}\mathbb{Z}/q_{1}\mathbb{Z}\times q_{2}^{\prime}\mathbb{Z}/q_{2}\mathbb{Z}.$ (3.15)
Proof.
Let $\mathbb{P}_{i},i=1,2$ be the projection from $\mathbb{Z}/q_{1}\mathbb{Z}\times\mathbb{Z}/q_{2}\mathbb{Z}$ to the $i$-th component. Without loss of generality we assume $q_{1}=q_{2}^{\alpha}$ for some $0<\alpha\leq 1$. Since $|A|>(q_{1}q_{2})^{1-\delta}$, there is $x_{0}\in A$ such that
|$|\{x\in A:\mathbb{P}_{1}(x)=x_{0}\}|>(q_{1}q_{2})^{1-\delta}/q_{1}>q_{2}^{1-\delta-\alpha\delta},$
so there is $A^{\prime}\subset A-A$, $|A^{\prime}|>q_{2}^{1-\delta-\alpha\delta}$, $\mathbb{P}_{1}(A^{\prime})=\{0\}$.
Similarly, there is $B^{\prime}\subset B-B$, $|B^{\prime}|>(q_{1}q_{2})^{1-\delta}/q_{1}>q_{2}^{1-\delta-\delta\alpha}$, $\mathbb{P}_{1}(B^{\prime})=0$. Applying Lemma 3.13, we obtain $q_{2}^{\prime}|q_{2},q_{2}^{\prime}<q_{2}^{\frac{12(\delta+\alpha\delta)}{5}}$ such that
$\sum_{48}(AB-AB)\supset\sum_{24}(A^{\prime}B^{\prime}-A^{\prime}B^{\prime})\supset\{0\}\times q_{2}^{\prime}\mathbb{Z}/q_{2}\mathbb{Z}.$ (3.16)
If $\alpha<5\delta$, then one can take $q_{1}^{\prime}=q_{1}$ and we have $q_{1}^{\prime}q_{2}^{\prime}<q_{2}^{10\delta}$, and (3.16) gives Lemma 3.13.
If $\alpha>5\delta$, then there exists $A^{\prime\prime}\subset A-A,B^{\prime\prime}\subset B-B$, such that $|A^{\prime\prime}|,|B^{\prime\prime}|>q_{1}^{1-\delta-\frac{\delta}{\alpha}},\mathbb{P}_{2}(A^{\prime\prime})=\mathbb{P}_{2}(B^{\prime\prime})=\{0\}$. The exponent $1-\delta-\delta/\alpha$ exceeds $3/4$, so applying Lemma 3.13, we obtain $q_{1}^{\prime}|q_{1}$, $q_{1}^{\prime}<q_{1}^{\frac{12(\delta+\frac{\delta}{\alpha})}{5}}$, such that
$\sum_{48}(AB-AB)\supset\sum_{24}(A^{\prime\prime}B^{\prime\prime}-A^{\prime\prime}B^{\prime\prime})\supset q_{1}^{\prime}\mathbb{Z}/q_{1}\mathbb{Z}\times\{0\}.$ (3.17)
Adding (3.16) and (3.17), we obtain Lemma 3.15, with
$q_{1}^{\prime}q_{2}^{\prime}<q_{1}^{\frac{12(\delta+\frac{\delta}{\alpha})}{5}}q_{2}^{\frac{12(\delta+\alpha\delta)}{5}}=(q_{1}q_{2})^{\frac{24\delta}{5}}.$
Proof of Proposition 3.12.
. Since $|A|>(q_{1}q_{2})^{3-\delta}$, by the pigeon hole principle, there exists $\vec{v}\in(\mathbb{Z}/q_{1}\mathbb{Z})^{2},\vec{w}\in(\mathbb{Z}/q_{2}\mathbb{Z})^{2}$, such that the cardinality of the set
$\{(\gamma_{1},\gamma_{2}):\gamma_{1}\text{ has lower row }\vec{v},\gamma_{2}\text{ has lower row }\vec{w}\}$
exceeds $(q_{1}q_{2})^{1-\delta}$. This implies the cardinality of the set
$A_{1}:=A\cdot A^{-1}\cap\left\{\left(\left(\begin{matrix}1&m\\ 0&1\end{matrix}\right),\left(\begin{matrix}1&n\\ 0&1\end{matrix}\right)\right):m\in\mathbb{Z}/q_{1}\mathbb{Z},n\in\mathbb{Z}/q_{2}\mathbb{Z}\right\}$
exceeds $(q_{1}q_{2})^{1-\delta}$.
Similarly, if we let
$A_{2}:=A\cdot A^{-1}\cap\left\{\left(\left(\begin{matrix}1&0\\ m&1\end{matrix}\right),\left(\begin{matrix}1&0\\ n&1\end{matrix}\right)\right):m\in\mathbb{Z}/q_{1}\mathbb{Z},n\in\mathbb{Z}/q_{2}\mathbb{Z}\right\}.$
Then, $|A_{2}|>(q_{1}q_{2})^{1-\delta}$.
Define an equivalence relation $\sim$ on $\text{SL}_{2}(\mathbb{Z}/q_{1}\mathbb{Z})\times\text{SL}_{2}(\mathbb{Z}/q_{2}\mathbb{Z})$ as:
$(\gamma_{1},\gamma_{2})\sim(\gamma_{1}^{\prime},\gamma_{2}^{\prime})$
if and only if the second rows of $\gamma_{1}$, $\gamma_{1}^{\prime}$ are the same up to a scaler in $(\mathbb{Z}/q_{1}\mathbb{Z})^{*}$, and the second rows of $\gamma_{2}$ $\gamma_{2}^{\prime}$ are the same up to a scaler in $(\mathbb{Z}/q_{1}\mathbb{Z})^{*}$. There are at most $q_{1}q_{2}$ many such classes. By pigeon hole, there exists one class which contains at least $(q_{1}q_{2})^{2-\delta}$ many elements from $A$. This implies the set
$H_{0}=A\cdot A^{-1}\cap\left\{\left(\left(\begin{matrix}\lambda_{1}&x\\ 0&\lambda_{1}^{-1}\end{matrix}\right),\left(\begin{matrix}\lambda_{2}&y\\ 0&\lambda_{2}^{-1}\end{matrix}\right)\right):\lambda_{1}\in(\mathbb{Z}/q_{1}\mathbb{Z})^{*},x\in\mathbb{Z}/q_{1}\mathbb{Z},\lambda_{2}\in(\mathbb{Z}/q_{2}\mathbb{Z})^{*},y\in\mathbb{Z}/q_{2}\mathbb{Z}\right\}$
has cardinality $>(q_{1}q_{2})^{2-\delta}$. By pigeon hole again, there exists $x_{0}\in\mathbb{Z}/q_{1}\mathbb{Z},y_{0}\in\mathbb{Z}/q_{2}\mathbb{Z}$ such that
$H=A\cdot A^{-1}\cap\left\{\left(\left(\begin{matrix}\lambda_{1}&x_{0}\\ 0&\lambda_{1}^{-1}\end{matrix}\right),\left(\begin{matrix}\lambda_{2}&y_{0}\\ 0&\lambda_{2}^{-1}\end{matrix}\right)\right):\lambda_{1}\in(\mathbb{Z}/q_{1}\mathbb{Z})^{*},\lambda_{2}\in(\mathbb{Z}/q_{2}\mathbb{Z})^{*}\right\}$
has cardinality $>(q_{1}q_{2})^{1-\delta}$.
We have the following elementary computation
$\left(\begin{matrix}\lambda&x_{0}\\ 0&\lambda^{-1}\end{matrix}\right)\left(\begin{matrix}1&x\\ 0&1\end{matrix}\right)\left(\begin{matrix}\lambda_{1}&x_{0}\\ 0&\lambda^{-1}\end{matrix}\right)^{-1}=\left(\begin{matrix}1&\lambda^{2}x\\ 0&1\end{matrix}\right).$ (3.18)
Using (3.18) and applying Lemma 3.14 to the set
$\left\{(\lambda_{1}^{2},\lambda_{2}^{2}):\left(\left(\begin{matrix}\lambda_{1}&x_{0}\\ 0&\lambda_{1}^{-1}\end{matrix}\right),\left(\begin{matrix}\lambda_{2}&y_{0}\\ 0&\lambda_{2}^{-1}\end{matrix}\right)\right)\in H\right\}$
and
$\left\{(m,n):\left(\left(\begin{matrix}1&m\\ 0&1\end{matrix}\right),\left(\begin{matrix}1&n\\ 0&1\end{matrix}\right)\right)\in A_{1}\right\}$
with exponent $1-2\delta$, we obtain $Q_{1}|q_{1},Q_{2}|q_{2}$, $Q_{1}Q_{2}<(q_{1}q_{2})^{20\delta}$ such that
$(A\cdot A^{-1})^{288}\supset\left\{\left(\left(\begin{matrix}1&Q_{1}\mathbb{Z}/q_{1}\mathbb{Z}\\ 0&1\end{matrix}\right),\left(\begin{matrix}1&Q_{2}\mathbb{Z}/q_{2}\mathbb{Z}\\ 0&1\end{matrix}\right)\right)\right\}$
Similarly, we can obtain $Q_{1}^{\prime}|q_{1},Q_{2}^{\prime}|q_{2}$, $Q_{1}^{\prime}Q_{2}^{\prime}<(q_{1}q_{2})^{20\delta}$ such that
$(A\cdot A^{-1})^{288}\supset\left\{\left(\left(\begin{matrix}1&0\\ Q_{1}^{\prime}\mathbb{Z}/q_{1}\mathbb{Z}&1\end{matrix}\right),\left(\begin{matrix}1&0\\ Q_{2}^{\prime}\mathbb{Z}/q_{2}\mathbb{Z}&1\end{matrix}\right)\right)\right\}$
Let $Q_{1}^{*}=\text{lcm}(Q_{1},Q_{1}^{\prime})$, $Q_{2}^{*}=\text{lcm}(Q_{2},Q_{2}^{\prime})$. It is an elementary exercise to check that for $m\leq{n}/{2}$, any element of the group
$\left\{\left(\begin{matrix}a&b\\ c&d\end{matrix}\right)\in\text{SL}_{2}(\mathbb{Z}/p^{n}\mathbb{Z}):a,d\equiv 1\ (\operatorname{mod}p^{2m}),b,c\equiv 0\ (\operatorname{mod}p^{2m})\right\}$
can be written as a $a_{1}b_{1}a_{2}b_{2}a_{3}$, where $a_{1},a_{2},a_{3}\in\left(\begin{matrix}1&p^{m}\mathbb{Z}/p^{n}\mathbb{Z}\\ 0&1\end{matrix}\right)$ and $b_{1},b_{2}\in\left(\begin{matrix}1&0\\ p^{m}\mathbb{Z}/p^{n}\mathbb{Z}&1\end{matrix}\right)$. From this it follows that if we let $q_{1}^{\prime}=\text{gcd}((Q_{1}^{*})^{2},q_{1})$, $q_{2}^{\prime}=\text{gcd}((Q_{2}^{*})^{2},q_{2})$, then $q_{1}^{\prime}q_{2}^{\prime}\leq q^{40\delta}$, and
$(A\cdot A^{-1})^{1440}\supset\mathrm{SL}_{2}(q_{1}^{\prime}\mathbb{Z}/q_{1}\mathbb{Z})\times\mathrm{SL}_{2}(q_{2}^{\prime}\mathbb{Z}/q_{2}\mathbb{Z}).$
Following the same method, we can prove
Lemma 3.19.
For any $0<\delta<\frac{1}{10}$, the following holds: Let $A\subset\mathrm{SL}_{2}(\mathbb{Z}/q\mathbb{Z})$ be symmetric and $|A|>q^{3-\delta}$. Then there exists $q^{\prime}|q,q^{\prime}<q^{10\delta}$ such that
$A^{720}\supset\Gamma(q^{\prime})/\Gamma(q).$
We need Lemma 3.19 to prove the following bounded generation result for $\text{ASL}_{2}$.
Proposition 3.20.
For any $0<\delta<10^{-4}$, the following holds: Let $q_{2}|q_{1}$ and let $A\subset\mathrm{SL}_{2}(\mathbb{Z}/q_{1}\mathbb{Z})\ltimes(\mathbb{Z}/q_{2}\mathbb{Z})^{2}$ be symmetric and $|A|>(q_{1}^{3}q_{2}^{2})^{1-\delta}$. Then there exists $q_{1}^{\prime}|q_{1},q_{2}^{\prime}|q_{2},q_{1}^{\prime}<q_{1}^{50\delta},q_{2}^{\prime}<q_{1}^{55\delta}$, such that
$A^{7204}\supset\Gamma(q_{1}^{\prime})/\Gamma(q_{1})\ltimes(q_{2}^{\prime}\mathbb{Z}/q_{2}\mathbb{Z})^{2}.$
Proof.
Define an equivalence relation on $A$:
$(\gamma_{1},v_{1})\sim(\gamma_{2},v_{2})\text{ if and only if }\gamma_{1}^{-1}v_{1}=\gamma_{2}^{-1}v_{2}.$
Since $|A|>(q_{1}^{3}q_{2}^{2})^{1-\delta}$ and the number of equivalence classes is bounded by $q_{2}^{2}$, there is one class $B$ such that $|B|>q_{1}^{3(1-\delta)}q_{2}^{-2\delta}>q_{1}^{3-5\delta}$. Take any $b\in B$, then $Bb^{-1}>q_{1}^{3-5\delta}$ and the unipotent part of $Bb^{-1}$ is 00. Applying Lemma 3.19, we have
$A^{1440}\supset(Bb^{-1})^{720}\supset\Gamma(q_{1}^{\prime})/\Gamma(q)\ltimes\{0\}$ (3.21)
for some $q_{1}^{\prime}<q_{1}^{50\delta}$.
By the pigeon hole principle, there exists $\gamma_{0}\in\mathrm{SL}_{2}(\mathbb{Z}/q_{1}\mathbb{Z})$ such that $B^{\prime}=\{(\gamma_{0},w),w\in\mathbb{Z}/q_{2}\mathbb{Z}\}\cap A$ has at least $\frac{|A|}{q_{1}^{3}}=q_{1}^{-3\delta}q_{2}^{2-2\delta}$. Take any $b^{\prime}\in B^{\prime}$. It follows that there exists $(1,w)\in B^{\prime}b^{\prime-1}$, $w=q^{\prime}w_{0}$, where $q^{\prime}|q_{2},q^{\prime}\leq q_{1}^{3\delta}q_{2}^{2\delta}$ and $w_{0}$ is a primitive vector in $(\mathbb{Z}/q_{2}\mathbb{Z})^{2}$.
It is an elementary exercise to check that for any primitive vector $v\in(\mathbb{Z}/p^{n}\mathbb{Z})^{2}$, any $m\leq n$,
$\Gamma(p^{m})/\Gamma(p^{n})v-\Gamma(p^{m})/\Gamma(p^{n})v\supset(p^{m}\mathbb{Z}/p^{n}\mathbb{Z})^{2}.$ (3.22)
Let $L\subset A^{2882}$ be the collection of conjugates of $(1,w_{0})$ by $\Gamma(q_{1}^{\prime})/\Gamma(q)\ltimes\{0\}$. Using (3.22), we obtain
$A^{5764}\supset LL^{-1}\supset(1,\text{gcd}(q_{1}^{\prime}q^{\prime},q_{2})\mathbb{Z}/q_{2}\mathbb{Z})$ (3.23)
Multiplying the RHS of (3.21) and the RHS of (3.23) we obtain Proposition 3.20 with $q_{2}^{\prime}=\text{gcd}(q_{1}^{\prime}q^{\prime},q_{2})$.
Let $u,v\in\mathbb{Z}$ and $f_{u,v}:\mathbb{Z}^{6}\rightarrow\mathbb{Z}^{2}$ be given by
$f_{u,v}(A,B,C,D,E,F)=\left(\left(\begin{matrix}A&B\\ C&D\end{matrix}\right)-1\right)\left(\begin{matrix}u\\ v\end{matrix}\right)-\left(\begin{matrix}E\\ F\end{matrix}\right).$
It is an elementary exercise to check
$H_{u,v}:=\left\{\left(\left(\begin{matrix}A&B\\ C&D\end{matrix}\right),\left(\begin{matrix}E\\ F\end{matrix}\right)\right)\in G\bigg{|}f_{u,v}=0\right\}.$
is a subgroup of $\text{SL}_{2}(\mathbb{Z})\ltimes\mathbb{Z}^{2}$.
The following lemma shows if a sectional map from (a large subgroup of) $\text{SL}_{2}(\mathbb{Z}/p^{n}\mathbb{Z})$ to $\text{SL}_{2}(\mathbb{Z}/p^{n}\mathbb{Z})\ltimes(\mathbb{Z}/p^{n}\mathbb{Z})^{2}$ is a homomorphism, then the image must be close to $H_{u,v}$.
Proposition 3.24.
Let $1\leq m<\frac{n}{4}$ and let $H$ be a subgroup of $\Gamma(p^{m})/\Gamma(p^{n})\ltimes(\mathbb{Z}/p^{n-m}\mathbb{Z})^{2}$ such that $\mathbb{P}_{0}(H)=\Gamma(p^{m})/\Gamma(p^{n})$ and $|H|=|\Gamma(p^{m})/\Gamma(p^{n})|$. Then there exist some $u,v\in\frac{1}{p^{m}}\mathbb{Z}$ such that
$H\cap\left(\Gamma(p^{2m})/\Gamma(p^{n})\ltimes(\mathbb{Z}/p^{n}\mathbb{Z})^{2}\right)(\operatorname{mod}p^{\lfloor\frac{n}{4}\rfloor})\subset H_{u,v}(\operatorname{mod}p^{\lfloor\frac{n}{4}\rfloor})$
Proof.
Lemma 3.24 follows from proofs of a series of claims.
Claim 1: Suppose $(g,w)\in H$. If $p^{k}|g-1$, then $p^{k-m}|w$.
Proof of Claim 1. We prove by backward induction on $k$. When $k=n$, we must have $w=0$ by assumptions on $H$. Now let $k<n$ and assume the statement is true for $k+1$. Write $g=1+p^{k}M$. Since for $t\in\mathbb{N}$, $g^{t}\equiv 1+tp^{k}M\pmod{p^{\min\{2k,n\}}}$, we have
$(g,w)^{p}=(g^{p},(1+g+g^{2}+\cdots+g^{p-1})w)$
with $g^{p}\in\Gamma(p^{k+1})/\Gamma(p^{n})$ and
$(1+g+g^{2}+\cdots+g^{p-1})w\equiv(g^{p},p(1+\frac{(p-1)}{2}p^{k}M)w)(\operatorname{mod}{p^{\min\{2k,n-m\}}}).$
By induction hypothesis,
$(1+\frac{p-1}{2}p^{k}M)pw\in(p^{k-m+1}\mathbb{Z}/p^{n-m}\mathbb{Z})^{2}.$
Since $1+\frac{p-1}{2}p^{k}M$ is non-singular $\operatorname{mod}p$, we get
$pw\in(p^{k-m+1}\mathbb{Z}/p^{n-m}\mathbb{Z})^{2}\Rightarrow w\in(p^{k-m}\mathbb{Z}/p^{n-m}\mathbb{Z})^{2}.$
This proves Claim 1.
Claim 2: If
$A_{1}=(g_{1},w_{1})=\left(\left(\begin{matrix}1&p^{k}\\ 0&1\end{matrix}\right),\left(\begin{matrix}p^{k-m}a_{1}\\ p^{k-m}b_{1}\end{matrix}\right)\right)\in H$
and
$A_{2}=(g_{2},w_{2})=\left(\left(\begin{matrix}1&0\\ p^{k}&1\end{matrix}\right),\left(\begin{matrix}p^{k-m}a_{2}\\ p^{k-m}b_{2}\end{matrix}\right)\right)\in H,$
with $m\leq k\leq\frac{n}{4}$, then $a_{1},b_{2}\equiv 0\pmod{p^{k}}$ if $p\neq 3$, and $a_{1},b_{2}\equiv 0\pmod{p^{k-1}}$ if $p=3$.
Proof of Claim 2. We compute
$A_{3}=A_{1}A_{2}A_{1}^{-1}A_{2}^{-1}\equiv\left(\left(\begin{matrix}1+p^{2k}&0\\ 0&1-p^{2k}\end{matrix}\right),p^{2k-m}\left(\begin{matrix}b_{2}\\ -a_{1}\end{matrix}\right)\right)(\operatorname{mod}p^{3k},p^{3k-m})$
and
$A_{4}=A_{3}A_{1}A_{3}^{-1}A_{1}^{-1}\equiv\left(\left(\begin{matrix}1&2p^{3k}\\ 0&1\end{matrix}\right),p^{3k-m}\left(\begin{matrix}2a_{1}\\ -b_{1}\end{matrix}\right)\right)(\operatorname{mod}p^{4k},p^{4k-m}).$
On the other hand,
$A_{5}=A_{1}^{2p^{2k}}\equiv\left(\left(\begin{matrix}1&2p^{3k}\\ 0&1\end{matrix}\right),p^{3k-m}\left(\begin{matrix}2a_{1}\\ 2b_{1}\end{matrix}\right)\right)(\operatorname{mod}p^{4k},p^{4k-m}).$
Applying Claim 1 to $A_{4}A_{5}^{-1}$, we obtain $3b_{1}\equiv 0(\operatorname{mod}p^{k})$. Similarly, $3a_{2}\equiv 0(\operatorname{mod}p^{k})$. Claim 2 thus follows.
Claim 3: If $41\leq m\leq n/$ 4, $\left(\left(\begin{matrix}1&p^{m}\\ 0&1\end{matrix}\right),\left(\begin{matrix}u_{1}\\ v_{1}\end{matrix}\right)\right)\in H$, then $v_{1}\equiv 0\pmod{p^{\lfloor\frac{n}{4}\rfloor}}$.
Proof of Claim 3. consider
$A_{6}=A_{1}^{p^{t}}=\left(\left(\begin{matrix}1&p^{m+t}\\ 0&1\end{matrix}\right),\left(\begin{matrix}\cdots\\ p^{t}v_{1}\end{matrix}\right)\right)\in H.$
If taking $t=\lfloor\frac{n}{4}\rfloor-m$, by Claim 2 we get $v_{1}\equiv 0\pmod{p^{\lfloor\frac{n}{4}\rfloor}}.$ This finishes the proof of Claim 3.
By the same consideration, we obtain
Claim 4: If $41\leq m\leq n/$ 4, $\left(\left(\begin{matrix}1&0\\ p^{m}&1\end{matrix}\right),\left(\begin{matrix}u_{2}\\ v_{2}\end{matrix}\right)\right)\in H$, then $u_{2}\equiv 0\pmod{p^{\lfloor\frac{n}{4}\rfloor}}$.
Proposition 3.24 thus follows by taking $u=\frac{1}{p^{m}}v_{2}$ and $v=\frac{1}{p^{m}}u_{1}$ since
$\left(\left(\begin{matrix}1&p^{m}\\ 0&1\end{matrix}\right),\left(\begin{matrix}u_{1}\\ \cdot\end{matrix}\right)\right),\left(\left(\begin{matrix}1&0\\ p^{m}&1\end{matrix}\right),\left(\begin{matrix}\cdot\\ v_{2}\end{matrix}\right)\right)\in H_{u,v}(\operatorname{mod}p^{\lfloor\frac{n}{4}\rfloor}),$
and that $\left(\begin{matrix}1&p^{m}\\ 0&1\end{matrix}\right)$ and $\left(\begin{matrix}1&0\\ p^{m}&1\end{matrix}\right)$ generate $\Gamma(p^{2m})/\Gamma(p^{n})$. ∎
4. Preliminaries on Random Walks
Recall that $S$ be a finitely symmetric set on ${\text{SL}}_{2}(\mathbb{Z})\times{\text{SL}}_{2}(\mathbb{Z})$ such that $\langle S\rangle$ is Zariski-dense, and $\chi_{S}$ be the uniform probability measure supported on $S$. The following lemmas are quantitative statements about non-concentration of self-convolutions of $\chi_{S}$ in proper subvarieties.
For the discussion in the following in this section, we prefix a primitive linear form
$L\left(\left(\begin{matrix}x_{1}&y_{1}\\ z_{1}&w_{1}\end{matrix}\right),\left(\begin{matrix}x_{2}&y_{2}\\ z_{2}&w_{2}\end{matrix}\right)\right)=X_{1}x_{1}+Y_{1}y_{1}+Z_{1}z_{1}+W_{1}w_{1}+X_{2}x_{2}+Y_{2}y_{2}+Z_{2}z_{2}+W_{2}w_{2},$
i.e., $\text{gcd}(X_{1},Y_{1},Z_{1},W_{1},X_{2},Y_{2},Z_{2},W_{2})=1$.
Proposition 4.1.
Let $S$ be a finite symmetric set on ${\text{SL}}_{2}(\mathbb{Z})\times{\text{SL}}_{2}(\mathbb{Z})$ such that $\langle S\rangle$ is Zariski-dense in $\text{SL}_{2}\times\text{SL}_{2}$. Let $\chi_{S}$ be the uniform probability measure supported on $S$. There are constants $10<c_{S},c<$ 1 depending only on $S$ such that for $Q\in\mathbb{Z}_{+}$, for any $l>{c_{S}}{\log Q}$ and $n\in\mathbb{Z}$, we have
$\chi_{S}^{(l)}\left(\left\{g\in{\text{SL}}_{2}(\mathbb{Z})\times{\text{SL}}_{2}(\mathbb{Z})|L(g)\equiv n(\operatorname{mod}Q)\right\}\right)<Q^{-c}.$
Proposition 4.2.
There are constants $10<c_{S},c<$ 1 depending only on $S$ such that the following holds. Let $Q\in\mathbb{Z}_{+}$ large enough and $\xi=(\xi_{1},\xi_{2}),\eta=(\eta_{1},\eta_{2})\in\text{Mat}_{2}(\mathbb{Z})\times\text{Mat}_{2}(\mathbb{Z})$ satisfy
$\text{Tr}(\xi_{1})=\text{Tr}(\xi_{2})=\text{Tr}(\eta_{1})=\text{Tr}(\eta_{2})=0$
and either for every $p|Q$, $\pi_{p}(\xi_{1}),\pi_{p}(\xi_{2})\neq 0$, or for every $p|Q$, $\pi_{p}(\eta_{1}),\pi_{p}(\eta_{2})\neq 0$. Then for $l>c_{S}\log Q$ and any $n\in\mathbb{Z}$,
$\chi_{S}^{(l)}(\{(g_{1},g_{2})\in{\text{SL}}_{2}(\mathbb{Z})\times{\text{SL}}_{2}(\mathbb{Z})|\text{Tr}(g_{1}\xi_{1}g_{1}^{-1}\eta_{1})+\text{Tr}(g_{2}\xi_{2}g_{2}^{-1}\eta_{2})\equiv n(\operatorname{mod}{Q})\})<Q^{-c}.$
In the $\mathrm{ASL}_{2}$ case, we also have the following analogue about non-concentration in proper subvarieties.
Proposition 4.4.
Let $S$ be a finite symmetric set on $\mathrm{SL}_{2}(\mathbb{Z})\ltimes\mathbb{Z}^{2}$ such that $\langle S\rangle$ is Zariski-dense in $\mathrm{SL}_{2}(\mathbb{Z})\ltimes\mathbb{Z}^{2}$. Let $\chi_{S}$ be the uniform probability measure supported on $S$. There are constants $10<c_{S},c<$ 1 depending only on $S$ such that for $Q\in\mathbb{Z}_{+}$, for any $l>{c_{S}}{\log Q}$ and $n\in\mathbb{Z}$, we have
$\chi_{S}^{(l)}\left(\left\{g\in\mathrm{SL}_{2}(\mathbb{Z})\ltimes\mathbb{Z}^{2}|L(g)\equiv n(\operatorname{mod}Q)\right\}\right)<Q^{-c},$
where $L$ is a primitive linear form on $\mathrm{SL}_{2}(\mathbb{Z})\ltimes\mathbb{Z}^{2}$.
We focus on proving Proposition 4.1; Proposition 4.2 and Proposition 4.4 follow in a similar way (see also Proposition 4.1 in [BG09]). We first show that Proposition 4.1 follow from
Lemma 4.5.
There are constants $10<c_{S},c_{0}<$ 1 depending only on $S$ such that for any $Q\in\mathbb{Z}_{+},n\in\mathbb{Z}$ and for $l\in\mathbb{Z}_{+}$ with $1\ll_{S}l<c_{S}\log Q$, we have
$\chi_{S}^{(l)}(\{g\in\text{SL}_{2}(\mathbb{Z})\times\text{SL}_{2}(\mathbb{Z})|L(g)\equiv n\pmod{{Q}}\})<e^{-c_{0}l}.$ (4.6)
Proof of Proposition 4.1 assuming Lemma 4.5.
Let $c_{S},c_{0}$ be the constants given by Lemma 4.5.
Let $l_{0}=[c_{S}\log Q]$ and write $\chi_{S}^{(l)}=\chi_{S}^{l_{0}}*\chi_{S}^{(l-l_{0})}$. For any $g^{\prime}$ in the support of $\chi_{S}^{l-l_{0}}$, let $L_{g^{\prime}}(g)=L(gg^{\prime})$. Clearly, $L_{g^{\prime}}$ is also primitive, so Lemma 4.5 is applicable to $L_{g^{\prime}}$. Therefore,
$\chi_{S}^{(l)}\left(\left\{g\in{\text{SL}}_{2}(\mathbb{Z})\times{\text{SL}}_{2}(\mathbb{Z})|L(g)\equiv n(\operatorname{mod}Q)\right\}\right)$
$=$ $\sum_{g^{\prime}\in\Lambda}\chi_{S}^{(l_{0})}\left(\left\{g\in\text{SL}_{2}(\mathbb{Z})\times\text{SL}_{2}(\mathbb{Z})|L_{(g^{\prime})^{-1}}(g)\equiv n(\operatorname{mod}Q)\right\}\right)\chi_{S}^{(l-l_{0})}(g^{\prime})$
$<$ $e^{-c_{0}l_{0}}<Q^{-c_{0}c_{S}}$
Proposition 4.1 is thus proved with $c=c_{0}c_{S}$
Now we focus on proving Lemma 4.5.
Lemma 4.7 (Non-concentration at Archimedean place).
There is a constant $c_{0}>0$ depending only on $S$ such that for any $l\gg_{S}1$, any $n\in\mathbb{Z}$, we have
$\chi_{S}^{(l)}\left(\left\{g\in\mathrm{SL}_{2}(\mathbb{Z})\times\mathrm{SL}_{2}(\mathbb{Z}):L(g)=n\right\}\right)<e^{-c_{0}l}.$
Proof.
By Theorem A , there exists an absolute constant $10<\lambda<$ 1 which is the upper bound for all eigenvalues of the family of the operators $T_{p}:l_{0}^{2}(\Lambda)\rightarrow l_{0}^{2}(\Lambda)$ defined at (2.1). So
$\|\chi_{S}^{(l)}-\dfrac{1}{|\pi_{p}(\Gamma)|}\boldsymbol{1}_{\pi_{p}(\Gamma)}\|_{2}\leq\lambda^{l}$ (4.8)
for all prime $p$. Then since $|\pi_{p}(\Gamma)|\approx p^{6}$, if $l>\frac{6\log p}{(\log 1/\lambda)}$,
$\chi_{S}^{(l)}(g)<\dfrac{2}{|\pi_{p}(\Gamma)|}$ (4.9)
for any $g\in\pi_{p}(\Gamma)$. Therefore,
$\chi_{S}^{(l)}\{g\in\mathrm{SL}_{2}(\mathbb{Z})\times{\text{SL}}_{2}(\mathbb{Z}):L(g)=n\}$
$\leq\chi_{S}^{(l)}\{g\in\mathrm{SL}_{2}(\mathbb{Z})\times{\text{SL}}_{2}(\mathbb{Z}):L(g)\equiv n(\operatorname{mod}{p})\}$
$<\frac{4}{p},$
by counting points in the subvariety of codimension $1$ determined by a primitive linear equation. So we are done by picking any prime $p\in[e^{\frac{l}{8}\log\frac{1}{\lambda}},e^{\frac{l}{7}\log\frac{1}{\lambda}}]$, which exists by the Prime Number Theorem when $l$ is sufficiently large. This finishes the proof of Lemma 4.7. ∎
We follow the method from Bourgain and Gamburd [BG09] to prove Lemma 4.5. The proof requires Effective Bézout Identity [BY91]. For readers’ convenience, we record it here:
Theorem C (The Effective Bézout Identity).
Let $\mathcal{P}_{1},\ldots,\mathcal{P}_{N}\in\mathbb{Z}[z_{1},\dots,z_{n}]$ without common zeros in $\mathbb{C}^{n}$ with $n\geq 2,\operatorname{deg}\mathcal{P}_{j}\leqslant d,d\geqslant 3$, $h\left(\mathcal{P}_{j}\right)\leqslant h$. Here $h(\mathcal{P})=\log\max\limits_{i}\left|a_{i}\right|$ for $\mathcal{P}(x)=a_{0}+a_{1}x+a_{2}x^{2}+\cdots+a_{n}x^{n}$ is the logarithm height function of polynomials. Then there is an integer $D\in\mathbb{Z}_{+}$, polynomials $\mathcal{Q}_{1},\ldots,\mathcal{Q}_{N}\in\mathbb{Z}[z_{1},\dots,z_{n}]$ such that
$\begin{gathered}\mathcal{P}_{1}\mathcal{Q}_{1}+\ldots+\mathcal{P}_{N}\mathcal{Q}_{N}=D,\\ \operatorname{deg}\mathcal{Q}_{j}\leq n(2n+1)d^{n},\\ h\left(\mathcal{Q}_{j}\right)\leqslant\mathfrak{X}(n)d^{8n+3}(h+\log N+d\log d),\\ \log D\leqslant\mathfrak{X}(n)d^{8n+3}(h+\log N+d\log d),\end{gathered}$
where $\mathfrak{X}(n)$ is an effective constant which only depends on $n$ and can be computed explicitly.
Proof of Lemma 4.5.
Write $Q=\prod_{i\in I}p_{i}^{n_{i}}$. We divide our proof into two cases $n=0$ and $n\neq 0$. We first deal with the case $n=0$.
By the primitivity of $L$, for each $p|Q$, at least one of
$t\in\mathcal{A}:=\{X_{1},Y_{1},Z_{1},W_{1},X_{2},Y_{2},Z_{2},W_{2}\}$
must be invertible $\operatorname{mod}\ p$. For each $t\in\mathcal{A}$, let
$Q_{t}=\prod_{\begin{subarray}{c}p_{i}^{n_{i}}||Q\\ \text{gcd}(p_{i},t)=1\end{subarray}}p_{i}^{n_{i}}$
Since $\prod_{t\in\mathcal{A}}Q_{t}\geq Q$, there exists $t\in\mathcal{A}$ such that
$Q^{\prime}:=Q_{t}\geq Q^{\frac{1}{8}}.$
Assume $t=X_{1}$ without loss of generality. Then $Q^{\prime}||Q$ and $(Q^{\prime},X_{1})=1$.
Now let
$\|(g_{1},g_{2})\|=\max\{\text{absolute values of coefficients of }g_{1}\text{ and }g_{2}\}.$
Let $C_{1}$ be an upper bound of $\|g\|$ for all $g\in\operatorname{supp}\chi_{S}$. Define
$\mathcal{G}=\left\{g\in\text{supp }[\chi_{S}^{(l)}]\mid L(g)\equiv 0\ (\operatorname{mod}{Q^{\prime}})\right\}.$
[…]
$\chi_{S}^{(l)}(\mathcal{G})<e^{-cl}.$
for some $c>0$. For each $\gamma=\left(\begin{pmatrix}x_{1}&y_{1}\\ z_{1}&w_{1}\end{pmatrix},\begin{pmatrix}x_{2}&y_{2}\\ z_{2}&w_{2}\end{pmatrix}\right)\in\mathcal{G}$ we introduce a linear polynomial $f_{\gamma}(\tilde{Y}_{1},\tilde{Z}_{1},\tilde{W}_{1},\tilde{X}_{2},\tilde{Y}_{2},\tilde{Z}_{2},\tilde{W}_{2})\in\mathbb{Q}[\tilde{Y}_{1},\tilde{Z}_{1},\tilde{W}_{1},\tilde{X}_{2},\tilde{Y}_{2},\tilde{Z}_{2},\tilde{W}_{2}]$ as follows:
$f_{\gamma}(\tilde{Y}_{1},\tilde{Z}_{1},\tilde{W}_{1},\tilde{X}_{2},\tilde{Y}_{2},\tilde{Z}_{2},\tilde{W}_{2})=x_{1}+\tilde{Y}_{1}y_{1}+\tilde{Z}_{1}z_{1}+\tilde{W}_{1}w_{1}+\tilde{X}_{2}x_{2}+{\tilde{Y}}_{2}y_{2}+\tilde{Z}_{2}z_{2}+\tilde{W}_{2}w_{2}$ (4.10)
Then we get
$X_{1}f_{\gamma}\left(Y_{1}\bar{X_{1}},Z_{1}\bar{X_{1}},W_{1}\bar{X_{1}},X_{2}\bar{X_{1}},Y_{2}\bar{X_{1}},Z_{2}\bar{X_{1}},W_{2}\bar{X_{1}}\right)\equiv 0\ (\operatorname{mod}{Q^{\prime}})\text{ for all }\gamma\in\mathcal{G}.$ (4.11)
Here $\bar{X_{1}}$ is the multiplicative inverse of $X_{1}$ mod $Q^{\prime}$.
Also, by definition of $\mathcal{G}$, the coefficients of $f_{\gamma}$, namely entries of $\gamma$, are bounded by $(2C_{1})^{l}$. Hence,
$h\left(f_{g}\right)<2l\log C_{1},$
We claim that there is a common zero $(\tilde{Y},\tilde{Z},\tilde{W},\tilde{A},\tilde{B},\tilde{C},\tilde{D})\in\mathbb{C}^{7}$ to the following system of equations:
$f_{\gamma}(Y,Z,W,A,B,C,D)=0\quad\text{ for all }\gamma\in\mathcal{G},$ (4.12)
[…]
Assume the claim fails to hold. We invoke Theorem C with $n=7,d=3,h=2l\log C_{1}$. It follows that there is an integer $M\in\mathbb{Z}_{+}$ and polynomials $\varphi_{l}$ of degree at most $b=7\times 15\times 3^{7}$ satisfying
$M=\sum_{l=1}^{N}F_{l}\varphi_{l}$ (4.13)
with
$0<\log M,h\left(\varphi_{l}\right)<\mathfrak{X}(7)3^{59}((\log C_{1}+\log(2|S|))l+3\log 3)<C_{2}l$ (4.14)
where $C_{2}=3^{61}\mathfrak{X}(7)\log 2C_{1}|S|$.
Now we take $(Y^{\prime},Z^{\prime},W^{\prime},A^{\prime},B^{\prime},C^{\prime},D^{\prime})\in\mathbb{Z}^{7}$ such that
$(Y^{\prime},Z^{\prime},W^{\prime},A^{\prime},B^{\prime},C^{\prime},D^{\prime})\equiv(Y_{1}\bar{X_{1}},Z_{1}\bar{X_{1}},W_{1}\bar{X_{1}},X_{2}\bar{X_{1}},Y_{2}\bar{X_{1}},Z_{2}\bar{X_{1}},W_{2}\bar{X_{1}})(\operatorname{mod}Q^{\prime}).$
It follows that
$M\equiv\sum_{l=1}^{N}F_{l}\varphi_{l}\Bigg{|}_{(Y^{\prime},Z^{\prime},W^{\prime},A^{\prime},B^{\prime},C^{\prime},D^{\prime})}\equiv 0\ (\operatorname{mod}{Q^{\prime}}),$
[…]
$x_{1}+\tilde{Y}y_{1}+\tilde{Z}z_{1}+\tilde{W}w_{1}+\tilde{A}x_{2}+\tilde{B}y_{2}+\tilde{C}z_{2}+\tilde{D}w_{2}=0$
Rescaling to get rid of the common denominator, we obtain $(\tilde{X}_{1},\tilde{Y}_{1},\tilde{Z}_{1},\tilde{W}_{1},\tilde{X}_{2},\tilde{Y}_{2},\tilde{Z}_{2},\tilde{W}_{2})\in\mathbb{Z}^{8}$ such that $\text{gcd}(\tilde{X}_{1},\tilde{Y}_{1},\tilde{Z}_{1},\tilde{W}_{1},\tilde{X}_{2},\tilde{Y}_{2},\tilde{Z}_{2},\tilde{W}_{2})=1$, and for all $\left(\left(\begin{matrix}x_{1}&y_{1}\\ z_{1}&w_{1}\end{matrix}\right)\left(\begin{matrix}x_{2}&y_{2}\\ z_{2}&w_{2}\end{matrix}\right)\right)\in\mathcal{G}$, we have
$\tilde{X}_{1}x_{1}+\tilde{Y}_{1}y_{1}+\tilde{Z}_{1}z_{1}+\tilde{W}_{1}w_{1}+\tilde{X}_{2}x_{2}+\tilde{Y}_{2}y_{2}+\tilde{Z}_{2}z_{2}+\tilde{W}_{2}w_{2}=0.$
We finish the proof of Lemma 4.5 of the case $n=0$ by applying Lemma 4.7 to the linear form $\tilde{L}$ determined by the constants $\tilde{X}_{1},\tilde{Y}_{1},\tilde{Z}_{1},\tilde{W}_{1},\tilde{X}_{2},\tilde{Y}_{2},\tilde{Z}_{2},\tilde{W}_{2}.$ For $1\ll_{S}l<c_{S}\log Q$,
$\chi_{S}^{(l)}\left(\left\{g\in\mathrm{SL}_{2}(\mathbb{Z})\times\mathrm{SL}_{2}(\mathbb{Z})\mid L(g)\equiv 0(\operatorname{mod}{Q})\right\}\right)$
[…]
=\ = $\chi_{S}^{(l)}(\mathcal{G})$
[…]
<\ < $e^{-c_{0}l}.$
where $c_{0}$ is the constant given in Lemma 4.7.
Now let us deal with the case $n\neq 0$. Let $C_{3}=3^{70}\mathfrak{X}(8)$, where $\mathfrak{X}(8)$ is the constant given by Theorem C, and take $c_{S}=\frac{1}{C_{3}}$. Write $Q_{1}=(Q,n)$.
Suppose $Q_{1}<Q^{\frac{1}{C_{3}}}$. First we can find $n^{\prime}=Q_{1}m\equiv 0(\operatorname{mod}Q)$ with $(m,Q)=1$. The equation $L(\gamma)=n(\operatorname{mod}Q)$ is the same as $L(\gamma)=Q_{1}m(\operatorname{mod}Q)$. Since $m$ is invertible mod $Q$, we can define
$f_{\gamma}(X,Y,Z,W,A,B,C,D)=xX+yY+zZ+wW+aA+bB+cD+dD-Q_{1}$
for $\gamma=\left(\left(\begin{matrix}x&y\\ z&w\end{matrix}\right),\left(\begin{matrix}a&b\\ c&d\end{matrix}\right)\right)$, and proceed as the $n=0$ case. The reason in the $n=0$ case we do not define $f_{\gamma}$ in this way is that we need to ensure a solution to $f_{\gamma}=0$ is nonzero. But if $n\neq 0$, $f_{\gamma}=0$ never admits a zero solution.
If $Q_{1}\geq Q^{\frac{1}{C_{3}}}$, then we can reduce this case to the $n=0$ case by considering the equation $L(g)\equiv 0(\operatorname{mod}Q_{1})$ instead of $L(g)\equiv n(\operatorname{mod}Q)$, and letting $c_{S}=\frac{1}{8C_{2}C_{3}}$. ∎
5. Bounded generation for simple factors
Let $\mathbb{P}_{0}$ be the projection from $\text{SL}_{2}(\mathbb{Z})\ltimes\mathbb{Z}^{2}$ to $\text{SL}_{2}(\mathbb{Z})$, and let $\mathbb{P}_{1}$ ($\mathbb{P}_{2}$, resp.) be the projection map from $\text{SL}_{2}(\mathbb{Z})\times\text{SL}_{2}(\mathbb{Z})$ to its first factor (second factor, resp.). The goal of this section is to prove, the following bounded generation result for simple factors of $\Lambda(\operatorname{mod}q)$, where $q$ is a product of large powers of primes:
Proposition 5.1.
Let $S,\epsilon$ be as given in Proposition 2.2. There are constants $c>0$ depending only on the generating set $S$ and $\epsilon$, and a constant $\rho=\rho(\delta)>0$, $C=C(\delta)\in\mathbb{Z}_{+}$, with $\rho=O_{S}(\delta)$, such that for any $q=\prod_{i\in I}p_{i}^{n_{i}},n_{i}>\frac{12}{\delta}$, if the conclusion of Proposition 2.2 fails, i.e.,
$|\pi_{q}(A\cdot A\cdot A)|\leq|\pi_{q}(A)|^{1+\delta},$ (5.2)
then we have
$\Gamma(q_{*}^{\{\rho\}})/\Gamma(q_{*})\subset\begin{cases}&(\mathbb{P}_{i}A)^{C}(\operatorname{mod}q_{*}),i=1,2,\hskip 8.53581pt\text{if }\hskip 5.69054pt\Lambda={\text{SL}}_{2}(\mathbb{Z})\times{\text{SL}}_{2}(\mathbb{Z}),\\ &\mathbb{(}P_{0}A)^{C}(\operatorname{mod}q_{*}),\hskip 45.5244pt\text{if }\hskip 5.69054pt\Lambda={\text{SL}}_{2}(\mathbb{Z})\ltimes\mathbb{Z}^{2}.\end{cases}$ (5.3)
where $q_{*}||q$ and $q_{*}\geq q^{c}$.
We focus on the case $\Lambda=\text{SL}_{2}(\mathbb{Z})\times\text{SL}_{2}(\mathbb{Z})$ and $i=1$. The proofs for the other two cases are identical. Our proof mostly follows the steps of [BG09].
5.1. A large set of commuting elements
We use Helfgott’s argument [Hel08] to create a large set of commuting elements under an appropriate modulus.
Let $c_{1},c_{2}$ be the implied constants for $c$ from Proposition 4.1 and Proposition 4.2 respectively. Let $A_{0}=A\cdot A\cap\Lambda(q_{0})$, where $q_{0}=\prod_{p|q_{l}}p$. Then
$\chi_{S}^{2l}(A_{0})>q^{-2\delta}q_{0}^{-12}>q^{-3\delta},$ (5.4)
where we have used the exponent for each prime divisor of $q$ exceeds $\frac{12}{\delta}$.
Let $\nu$ be the push forward of $\chi_{S}$ under the projection map $\mathbb{P}_{1}$ and let
$A_{1}=\mathbb{P}_{1}(A_{0}).$
For each $Q|q$, let
$\mathcal{E}_{1}(Q)=\{\gamma\in\text{SL}_{2}(\mathbb{Z}):\text{tr}(\gamma)^{2}-4\equiv 0(\operatorname{mod}Q))\},$
$\mathcal{E}_{2}(Q)=\left\{\gamma\in\text{SL}_{2}(\mathbb{Z}):\gamma_{21}\equiv 0(\operatorname{mod}Q))\right\}.$
Then Proposition 4.1 gives
$\nu^{(2l)}(\mathcal{E}_{1}(Q)),\nu^{(2l)}(\mathcal{E}_{2}(Q))<Q^{-c_{1}}.$
Take $\rho_{1}>0$, and let
$A_{2}=A_{1}-\cup_{\begin{subarray}{c}Q|q\\ Q\geq q^{\{\rho_{1}\}}\end{subarray}}(\mathcal{E}_{1}(Q)\cup\mathcal{E}_{2}(Q)),$
We should think of $\rho_{1}$ as a relative level depending only on $S$, so as $\rho_{2}$ in the following; the exact value of $\rho_{1}$ is determined at (5.58). Since the number of divisors of $q$ is bounded by $q^{0+}$, it follows that
$\nu^{(2l)}(A_{2})>q^{-3\delta-},$
if we let
$\delta<\frac{\rho_{1}c_{1}}{4}.$ (5.5)
A simple argument by contradiction shows that for each $\gamma\in A_{2}$, there exists $q_{\gamma}||q$, $q_{\gamma}>q^{\frac{1}{2}}$, such that for any $p^{n}||q_{\gamma}$,
$\text{tr}(\gamma)^{2}-4\neq 0(\operatorname{mod}p^{[4\rho_{1}n]}),$ (5.6)
$\gamma_{21}\neq 0(\operatorname{mod}p^{[4\rho_{1}n]}).$ (5.7)
Since again the number of divisors of $q$ is bounded by $q^{0+}$, there exists $q_{1}||q,q_{1}>q^{\frac{1}{2}}$ and a set $A_{3}\subset A_{2}$, with
$\nu^{(2l)}(A_{3})>q^{-3\delta-},$ (5.8)
[…]
Let
$Q_{1}=q_{1}^{\{\rho_{2}\}},$
where we will take $\rho_{2}\gg\rho_{1}$.
Let
$W(n):=\{\gamma\in\Gamma:\text{tr}(\gamma)=n(\operatorname{mod}Q_{1})\}.$
By Proposition 4.1, for any $n\in\mathbb{Z}$,
$\nu^{2l}(W(n))<Q_{1}^{-c_{1}}.$ (5.9)
If we take
${\delta}<{\frac{\rho_{2}c_{1}}{12}},$ (5.10)
then $\nu^{(2l)}(A_{3})>Q_{1}^{-\frac{c_{1}}{2}}$, and together with (5.9), it implies that there exists a set $A_{4}\subset A_{3}$, such that
$|A_{4}(\operatorname{mod}Q_{1})|>Q_{1}^{c_{1}/2}$ (5.11)
$\forall\gamma_{1},\gamma_{2}\in A_{4},\gamma_{1}\neq\gamma_{2}\Rightarrow\text{tr}(\gamma_{1})\neq\text{tr}(\gamma_{2})(\text{mod }Q_{1})$ (5.12)
Next we need to use our assumption (5.2) to get a control over the growth of $\pi_{Q_{1}}(A_{1})$. By Lemma 2.2 from [Hel08], (5.2) implies for any $l\geq 3$,
$|\pi_{q}(\prod_{l}A)|\leq\left(\frac{|\pi_{q}(A\cdot A\cdot A)|}{|\pi_{q}(A)|}\right)^{l-2}|\pi_{q}(A)|$ (5.13)
$<$ $|\pi_{q}(A)|^{1+\delta(l-2)}.$ (5.14)
It follows from Proposition 4.1 that for any $\gamma_{0}\in\Lambda_{q}$,
$\nu^{(2l)}\{\gamma\in\Lambda:\gamma\equiv\gamma_{0}(\operatorname{mod}Q_{1})\}<Q_{1}^{-c_{1}}.$ (5.15)
From (5.15) and (5.10), we conclude that
$|\pi_{Q_{1}}(A_{1})|>Q_{1}^{c_{1}}q^{-3\delta}>Q_{1}^{\frac{c_{1}}{2}}.$
Next we show small expansion for $A$ implies small expansion for $A_{1}$:
$|\pi_{Q_{1}}(A_{1}\cdot A_{1}\cdot A_{1})|<|\pi_{Q_{1}}(A_{1})|^{1+\frac{4\delta}{\rho_{2}c_{1}}}.$ (5.16)
To see this, suppose not, then
$|\pi_{q}(A^{4})|\geq$ $|\pi_{Q_{1}}(A_{1}\cdot A_{1}\cdot A_{1})|\cdot\max_{\gamma_{0}\in\Lambda_{Q_{1}}}|\pi_{Q_{1}}(A)\cap\{(\gamma_{0},\gamma):\gamma\in\text{SL}_{2}(\mathbb{Z}/q\mathbb{Z})\}|$
$\geq$ $|\pi_{Q_{1}}(A_{1})|^{1+\frac{4\delta}{\rho_{2}c_{1}}}\cdot\max_{\gamma_{0}\in\Lambda_{Q_{1}}}|\pi_{Q_{1}}(A)\cap\{(\gamma_{0},\gamma):\gamma\in\text{SL}_{2}(\mathbb{Z}/q\mathbb{Z})\}|$
$\geq$ $|\pi_{Q_{1}}(A_{1})|^{\frac{4\delta}{\rho_{2}c_{1}}}\cdot|\pi_{q}(A)|$
$>$ $|\pi_{q}(A)|^{1+\delta}$ (5.17)
which contradicts (5.2).
For each $\alpha\in A_{4}$, let $C_{\alpha}=\{\gamma\alpha\gamma^{-1}:\gamma\in A_{1}\}$. Since distinct elements in $A_{4}$ have distinct eigenvalues mod $Q_{1}$, the sets $\{\pi_{Q_{1}}(C_{\alpha})\}_{\alpha\in A_{1}}$ are mutually disjoint. Therefore,
$\sum_{\alpha\in A_{4}}|\pi_{Q_{1}}(C_{\alpha})|=|\cup_{\alpha\in A_{4}}\pi_{Q_{1}}(C_{\alpha})|<|\pi_{Q_{1}}(A_{1}^{3})|<|\pi_{Q_{1}}(A_{1})|^{1+\frac{4\delta}{\rho_{2}c_{1}}}.$
Thus, for some $\alpha_{0}=\left(\begin{matrix}\alpha_{11}&\alpha_{12}\\ \alpha_{21}&\alpha_{22}\end{matrix}\right)\in A_{4}$,
$|\pi_{Q_{1}}(C_{\alpha_{0}})|\leq|\pi_{Q_{1}}(A)|^{1+\frac{4\delta}{\rho_{2}c_{1}}}Q_{1}^{-c_{1}/2}<|\pi_{Q_{1}}(A)|Q_{1}^{-\frac{c_{1}}{4}}.$ (5.18)
provided that
$\delta<\frac{\rho_{2}c_{1}^{2}}{48}.$ (5.19)
To proceed, we first diagonalise $\alpha_{0}$ in an extended number field, i.e., let
$\lambda_{1}=\frac{\text{tr}(\alpha_{0})+\sqrt{\text{tr}(\alpha_{0})^{2}-4}}{2},\hskip 14.22636pt\lambda_{2}=\frac{\text{tr}(\alpha_{0})-\sqrt{\text{tr}(\alpha_{0})^{2}-4}}{2}$
be the two eigenvalues of $\alpha_{0}$. Let $K=\mathbb{Q}[\lambda_{1}]$, and so $\mathcal{O}=\mathcal{O}_{K}=\mathbb{Z}[\lambda_{1}]$. Let $M=\left(\begin{matrix}\lambda_{1}-\alpha_{22}&\lambda_{2}-\alpha_{22}\\ \alpha_{21}&\alpha_{21}\end{matrix}\right)$. Then
$M^{-1}\alpha_{0}M=\left(\begin{matrix}\lambda_{1}&0\\ 0&\lambda_{2}\end{matrix}\right).$
Since $\alpha_{0}$ satisfies the condition (5.6) for $\gamma$,
$\lambda_{1}\not\equiv\lambda_{2}(\operatorname{mod}p^{[4\rho_{1}n]})$ (5.20)
for any $p^{n}||q_{1}$.
From (5.18), there exists $x_{0}\in A_{4}$ and
$A_{5}:=\{x\in A:x\alpha_{0}x^{-1}=x_{0}\alpha_{0}x_{0}^{-1}(\operatorname{mod}Q_{1})\},$ (5.21)
such that
$|\pi_{Q_{1}}(A_{5})|>Q_{1}^{\frac{c_{1}}{4}}.$ (5.22)
Therefore, $A_{5}^{-1}\cdot x_{0}$ commute with $\alpha_{0}(\operatorname{mod}Q_{1})$ and
$|\pi_{Q_{1}}(A_{5}^{-1}\cdot x_{0})|>Q_{1}^{\frac{c_{1}}{4}}.$ (5.23)
Let $C_{1},C_{2},C_{3}$ be the implied constants of Theorem B for $\delta=\frac{c_{1}}{17}$.
The condition (5.23) implies that there exists $\upsilon\in[\frac{c_{1}}{24},\frac{1}{2}]$, such that
$|\pi_{Q_{1}^{\{\upsilon(1+\frac{1}{C_{1}})\}}}(A_{5})|>|\pi_{Q_{1}^{\{\upsilon\}}}(A_{5})|Q_{1}^{\frac{\upsilon c_{1}}{8C_{1}}}.$ (5.24)
If we let $A_{6}=\left((A_{5}^{-1}\cdot x_{0})\cdot(A_{5}^{-1}\cdot x_{0})^{-1}\right)\cap\Gamma(Q_{1}^{\{\upsilon\}})=A_{5}^{-1}A_{5}\cap\Gamma(Q_{1}^{\{\upsilon\}})$, then
$|\pi_{Q_{1}^{\{\upsilon(1+\frac{1}{C_{1}})\}}}(A_{6})|>Q_{1}^{\frac{\upsilon c_{1}}{8C_{1}}}.$ (5.25)
Let
$Q_{2}:=q_{1}^{\{\rho_{2}\upsilon-9\rho_{1}\}},$
where we will take $\rho_{2}\upsilon\gg 9\rho_{1}$. Since $\text{Det}(M)=\alpha_{21}\sqrt{\text{tr}(\alpha_{0})^{2}-4}$, from (5.7), (5.6), we have
$p^{[6\rho_{1}n]}\nmid\text{Det}(M),\forall p^{n}||q_{1}.$ (5.26)
Combining (5.26) with (5.20), we conclude that the matrix $M$ conjugate every matrix $\sigma$ in $A_{6}$ to a matrix in $\text{SL}_{2}(\mathcal{O})$ which is congruent to identity mod $Q_{2}$ and diagonal mod $Q_{2}^{\left\{1+\frac{1}{C_{1}}\right\}}$. In other words, there exists $a_{\sigma}\in\mathcal{O}_{K}$ and $a_{\sigma}\equiv 1(\operatorname{mod}Q_{2})$ such that
$M^{-1}\sigma M\equiv\left(\begin{matrix}a_{\sigma}&0\\ 0&a_{\sigma}^{-1}\end{matrix}\right)(\operatorname{mod}Q_{2}^{\left\{1+\frac{1}{C_{1}}\right\}}).$ (5.27)
It also follows from (5.25) that
$\Big{|}\pi_{Q_{2}^{\left\{1+\frac{1}{C_{1}}\right\}}}(A_{6})\Big{|}>Q_{2}^{\frac{c_{1}}{16C_{1}}}.$ (5.28)
if
$\rho_{1}<\frac{\upsilon c_{1}}{2^{5}3^{3}}\rho_{2},$ (5.29)
which is guaranteed if
$\rho_{1}<\frac{c_{1}^{2}}{2^{8}3^{4}}\rho_{2}.$ (5.30)
For each $\gamma\in A_{6}$, let ${q}_{\gamma}$ be the largest divisor of $Q_{2}^{\left\{1+\frac{1}{C_{1}}\right\}}$ such that $\gamma\equiv 1(\operatorname{mod}q_{\gamma})$. Obviously, $Q_{2}|q_{\gamma}$. Then we can write
$A_{6}=\sqcup_{Q_{2}|q^{\prime}|Q_{2}^{\left\{1+\frac{1}{C_{1}}\right\}}}B_{q^{\prime}},$
where $B_{q^{\prime}}=\{\gamma\in A_{6}:q_{\gamma}=q^{\prime}\}$. Since the number of divisors of $Q_{2}^{\left\{1+\frac{1}{C_{1}}\right\}}$ is bounded by $q^{0+}$, there exists $\tilde{q}$ such that $Q_{2}|\tilde{q}|Q_{2}^{\left\{1+\frac{1}{C_{1}}\right\}}$ and
$\Big{|}B_{\tilde{q}}(\operatorname{mod}Q_{2}^{\left\{1+\frac{1}{C_{1}}\right\}})\Big{|}>Q_{2}^{\frac{c_{1}}{16C_{1}}-}.$ (5.31)
5.2. Applying sum-product
Let $H=\{h_{\sigma}=(a_{\sigma}-a_{\sigma}^{-1})/\tilde{q}:\sigma\in B_{\tilde{q}}\}\subset\mathcal{O}$, where $a_{\sigma}$ is as given in (5.27). Since the map $a_{\sigma}\mapsto a_{\sigma}-a_{\sigma}^{-1}$ is bounded to one, we have
$\Big{|}H(\operatorname{mod}{Q_{2}^{\{\frac{1}{C_{1}}\}}})\Big{|}>Q_{2}^{\frac{c_{1}}{16C_{1}}-}.$ (5.32)
Recall $\mathcal{O}$ is a degree 2 extension of $\mathbb{Z}$. We apply Theorem B to the set $H$ with $\mathfrak{a}=(Q_{2}^{\{\frac{1}{C_{1}}\}})$ and $\delta=\frac{c_{1}}{17}$. Theorem B then produces some $\xi\in\mathcal{O}$, $\mathfrak{a}^{\prime}\supset\mathfrak{a}^{C_{1}}$, $c_{3}=c_{3}(c_{1})>0$ such that
$\sum_{C_{3}}H^{C_{2}}\supset\mathbb{Z}\xi(\operatorname{mod}\mathfrak{a}^{\prime}),$ (5.33)
and
$|\mathbb{Z}\xi(\operatorname{mod}\mathfrak{a}^{\prime})|>N(\mathfrak{a})^{c_{3}}>q_{1}^{\frac{\rho_{2}\upsilon c_{3}}{C_{1}}}>q_{1}^{\frac{\rho_{2}c_{1}c_{3}}{24C_{1}}}.$ (5.34)
We pass (5.34) to a local property. From (5.34), it follows that there exists $q_{2}||q_{1}$, $q_{2}>q_{1}^{\frac{c_{3}}{2C_{1}}}$, such that for any $p^{n}||q_{2}$, letting $\overline{\mathfrak{a}_{p}}=\langle\mathfrak{a}^{\prime},p^{n}\rangle$, the ideal of $\mathcal{O}$ generated by $\mathfrak{a}^{\prime}$ and $p^{n}$, we have
$|\mathbb{Z}\xi(\operatorname{mod}\overline{\mathfrak{a}_{p}})|>p^{\frac{\rho_{2}c_{1}c_{3}n}{48C_{1}}}.$
Since there are at most two prime ideals in $\mathcal{O}$ lying above $p$, there exists a prime ideal $\mathcal{P}_{p}\supset(p)$ in $\mathcal{O}$, $\mathcal{P}_{p}^{m_{p}}||\bar{\mathfrak{a}_{p}}$, such that
$|\mathbb{Z}\xi(\operatorname{mod}\mathcal{P}_{p}^{m_{p}})|>p^{\frac{\rho_{2}c_{1}c_{3}n}{96C_{1}}}.$ (5.35)
For simplicity, we assume for each $p|q_{3}$, the arising prime ideal $\mathcal{P}_{p}$ is unramified, so $v_{p}(\tilde{q})=v_{\mathcal{P}}(\tilde{q})$. The general case adds no additional technical challenge.
We obtain from (5.35), $\mathfrak{a}^{\prime}\supset\mathfrak{a}^{C_{1}}$ and the definition of $\mathfrak{a}$ that
$\frac{\rho_{2}c_{1}c_{3}}{96C_{1}}n<m_{p}<\rho_{2}n.$ (5.36)
It also follows from (5.35) that for any $p|q_{2}$,
$v_{p}(Q_{2})\leq v_{p}(\tilde{q})<v_{p}(Q_{2}^{\{1+\frac{1}{C_{1}}\}}).$ (5.37)
5.3. Constructing an appropriate element $\zeta$
We also need to find an element $\zeta=\left(\begin{matrix}a&b\\ c&d\end{matrix}\right)$ from $A_{1}$ which is very non-commutative with $\alpha_{0}=\left(\begin{matrix}\alpha_{11}&\alpha_{12}\\ \alpha_{21}&\alpha_{22}\end{matrix}\right)$. This leads us to work with the linear form
$\mathcal{L}_{1}\left(\left(\begin{matrix}a&b\\ c&d\end{matrix}\right)\right)=\alpha_{21}a+(\alpha_{22}-\alpha_{11})c-\alpha_{21}d,$
which gives the 2-1 entry of $\alpha_{0}\zeta-\zeta\alpha_{0}$. If $\mathcal{L}_{1}(\zeta)$ is non-zero, this certainly implies $\zeta$ and $\alpha_{0}$ are not commutative.
Let
$\mathcal{E}_{3}(Q)=\{\gamma\in\text{SL}_{2}(\mathbb{Z}):\mathcal{L}_{1}(\gamma)\equiv 0(\operatorname{mod}Q)\}.$
By Proposition 4.1, for any $Q|q_{2}$ ,
$\nu^{(2l)}(\mathcal{E}_{3}(Q))<Q^{-c_{1}},$ (5.38)
Define
$A_{7}=A_{1}-\cup_{\begin{subarray}{c}Q|q_{2}\\ Q>q_{2}^{\rho_{1}}\end{subarray}}\mathcal{E}_{3}(Q).$
By (5.4), $\nu^{(2l)}(A_{1})>q^{-3\delta}$. We also have $q_{2}>q_{1}^{\frac{c_{3}}{2C_{1}}}>q^{\frac{c_{3}}{4C_{1}}}$. Then
$\nu^{(2l)}(A_{7})>q^{-4\delta}$
if $q_{2}^{-\rho_{1}c_{1}}<q^{-4\delta}$, or
${\delta}<\frac{c_{1}c_{3}\rho_{1}}{16C_{1}}.$ (5.39)
Then we produce some $q_{3}||q_{2}$, $q_{3}>q_{2}^{\frac{1}{2}}$, $\zeta_{1}\in A_{1}$ such that for any $p^{n}||q_{3}$,
$L({\zeta_{1}})\neq 0(\operatorname{mod}p^{[2\rho_{1}n]}).$ (5.40)
From (5.40) and (5.20), we have $M\zeta_{1}M^{-1}=\left(\begin{matrix}a_{11}&a_{12}\\ a_{21}&a_{22}\end{matrix}\right)$ with $a_{12}$ or $a_{21}\not\equiv 0(\operatorname{mod}p^{[6n\rho_{1}]})$. Take any element $\zeta_{2}\in B_{\tilde{q}}$. The second inequality of (5.37) ensures that $v_{p}(\zeta_{2})=v_{p}(\tilde{q})$.
If we let $\zeta=\zeta_{1}\zeta_{2}\zeta_{1}^{-1}$ and write $M{\zeta}M^{-1}=\left(\begin{matrix}\zeta_{11},\zeta_{12}\\ \zeta_{21},\zeta_{22}\end{matrix}\right)$. Then
$v_{p}(M\zeta M^{-1}-1)\geq v_{p}(\tilde{q})-[6\rho_{1}n]$ (5.41)
and
$\min\{v_{p}(\zeta_{12}),v_{p}(\zeta_{21})\}\leq v_{p}(\tilde{q})+[10\rho_{1}n].$ (5.42)
5.4. Creating a line
Let
$\left\{\vec{e}=\left(\begin{matrix}0&1\\ 0&0\end{matrix}\right),\vec{f}=\left(\begin{matrix}0&0\\ 1&0\end{matrix}\right),\vec{h}=\left(\begin{matrix}1&0\\ 0&-1\end{matrix}\right)\right\}$
[…]
To proceed, we recall the following well known identity.
Lemma 5.43.
Let $m,m^{\prime}\in\mathbb{Z}_{+}$, and let $x,y\in\text{SL}_{2}(\mathcal{O})$, $x\equiv 1(\operatorname{mod}\mathcal{P}^{m}),y\equiv 1(\operatorname{mod}\mathcal{P}^{m^{\prime}})$. Then
$xyx^{-1}y^{-1}\equiv 1(\operatorname{mod}\mathcal{P}^{m+m^{\prime}}),$
and
$xyx^{-1}y^{-1}\equiv 1+xy-yx(\operatorname{mod}\mathcal{P}^{m+m^{\prime}+\min\{m,m^{\prime}\}})$
Lemma 5.43 has an obvious generalization to a general ideal by multiplicativity.
To apply Lemma 5.43, we take $x=M^{-1}\sigma M$ given at (5.27) and $y=M^{-1}\zeta M$, with corresponding moduli $\mathfrak{q}_{1}=\prod_{p|q_{3}}\mathcal{P}_{p}^{v_{\mathcal{P}}(\tilde{q})}$ and $\mathfrak{q}_{2}=\prod_{p|q_{3}}\mathcal{P}_{p}^{v_{\mathcal{P}}(\tilde{q})-[6\rho_{1}n]}$.
Then
$M^{-1}\sigma\zeta\sigma^{-1}\zeta^{-1}M\equiv 1(\operatorname{mod}\mathfrak{q}_{1}\mathfrak{q}_{2})$ (5.44)
and
$M^{-1}\sigma\zeta\sigma^{-1}\zeta^{-1}M\equiv 1+\tilde{q}h_{\sigma}\left(\zeta_{21}\vec{e}-\zeta_{12}\vec{f}\right)(\operatorname{mod}\mathfrak{q}_{1}\mathfrak{q}_{2}^{2})$ (5.45)
Applying Lemma 5.43 iteratively, for any $\sigma_{1},\sigma_{2},\cdots,\sigma_{C_{3}}\in A_{6}$, we have
$M^{-1}[\sigma_{C_{3}},\cdots[\sigma_{2},[\sigma_{1},\zeta]]\cdots]M\equiv 1(\operatorname{mod}(\mathfrak{q}_{1}^{C_{3}-1}\mathfrak{q}_{2}^{2})$ (5.46)
and
$M^{-1}[\sigma_{C_{3}},\cdots[\sigma_{2},[\sigma_{1},\zeta]]\cdots]M\equiv 1+\tilde{q}^{C_{3}}\left(\prod_{i=1}^{C_{3}}h_{\sigma_{i}}\right)X(\operatorname{mod}(\mathfrak{q}_{1}^{C_{3}}\mathfrak{q}_{2}^{2})$ (5.47)
where $X=\left(\begin{matrix}0&\zeta_{12}\\ (-1)^{C_{3}}\zeta_{21}&0\end{matrix}\right)$.
The sum structure is clear. For any $\sigma_{1,1},\cdots,\sigma_{C_{2},C_{3}}\in A_{6}$, we have
$(M^{-1}[\sigma_{1,C_{3}},\cdots[\sigma_{1,2},[\sigma_{1,1},\zeta]]M)\cdots(M^{-1}[\sigma_{C_{2},C_{3}},\cdots[\sigma_{C_{2,2}},[\sigma_{C_{{2},1}},\zeta]]\cdots]M)$
$\equiv$ $1+\tilde{q}^{C_{3}}\sum_{i=1}^{C_{2}}\left(\prod_{j=1}^{C_{3}}h_{\sigma_{i,j}}\right)X(\operatorname{mod}(\mathfrak{q}_{1}^{C_{3}}\mathfrak{q}_{2}^{2})$ (5.48)
From (5.33) and (5.4), we conclude there is a constant $K_{1}$ depending only on $c_{1}$, such that
$1+\tilde{q}^{C_{3}}\mathbb{Z}\xi X\subset M^{-1}A_{1}^{K_{1}}M(\operatorname{mod}\mathfrak{q}_{1}^{C_{3}}\mathfrak{q}_{2}\langle\mathfrak{a}^{\prime},\mathfrak{q}_{2}\rangle)$ (5.49)
Conjugating (5.49) back by $M$, noticing that $A_{1}^{K_{1}}$ is real and also having (5.35), (5.42), and (5.49) in mind, if we take
$\rho_{1}<\frac{c_{1}c_{3}}{28*192C_{1}}\rho_{2},$ (5.50)
then we obtain
Lemma 5.51.
There exists $q_{3}||q_{1}$, $q_{3}>q^{\frac{c_{1}}{8C_{1}}}$, $q_{3}^{\prime}|q_{3}^{\prime\prime}|q_{3}$, some $K_{1}=K_{1}(c_{1})>0$, a matrix $X_{0}\in{{\text{Lie(SL}}}_{2})(\mathbb{Z})$ coprime to $q_{3}$, such that
$1+q_{3}^{\prime}\mathbb{Z}X_{0}\subset A_{1}^{K_{1}}(\operatorname{mod}q_{3}^{\prime\prime})$ (5.52)
where for each $p^{n}||q_{3}$, $v_{p}(q_{3}^{\prime\prime})=O_{c_{1}}(\rho_{2}n)$ and $\frac{c_{1}c_{3}}{192C_{1}}\rho_{2}n<v_{p}(q_{3}^{\prime\prime}/q_{3}^{\prime})<{v_{p}(q_{3}^{\prime\prime})}/2$.
5.5. Creating a segment
Next, we want to find $g_{1},g_{2}$ from a bounded product of $A_{1}$ so that $g_{1},g_{2}$ conjugates $X_{0}$ to different directions.
[…]
If $\Lambda=\text{SL}_{2}(\mathbb{Z})\times\text{SL}_{2}(\mathbb{Z})$, apply Proposition 4.2 with $\xi_{1}=X_{0},\xi_{2}=\eta_{2}=0$, and a proper choice of $\eta_{1}$ so that $\text{Tr}(g\xi_{1}g^{-1}\eta_{1})=L_{1}(g)$. If $\Lambda=\text{SL}_{2}(\mathbb{Z})\ltimes\mathbb{Z}^{2}$, apply Proposition 4.4 with $L=L_{1}$. Let $c_{2}$ be the implied constant from Propositions 4.2 and 4.4. Then we obtain, for any $Q>q_{3}^{\rho_{1}}$, we have
$\nu^{(l)}(\{g\in\text{SL}_{2}(\mathbb{Z})|L_{1}\circ T_{1}(g\tilde{X}g^{-1})\equiv 0(\operatorname{mod}{Q})\})<Q^{-c_{2}},$
where $l$ is as given in Proposition 2.2. This implies, arguing the same as for the construction of $A_{3}$, that there is $q_{4}||q_{3}$, $q_{4}>(q_{3})^{\frac{1}{2}}$, and $g_{1}\in A_{1}$ such that for any $p^{n}||q_{5}$, $L_{1}\circ T_{1}(g_{1}X_{0}g_{1}^{-1})\neq 0(p^{[2\rho_{1}n]})$, if we take
$\delta<\frac{c_{1}c_{2}\rho_{1}}{48C_{1}}.$ (5.53)
Let $X_{1}=g_{1}X_{0}g_{1}^{-1}$, then $X_{1}\equiv a(1,0,0)+b\vec{v}(\operatorname{mod}q_{5})$, where $a,b\in\mathbb{Z}$, and $b$ satisfies, for any $p^{n}||q_{5}$, $p^{[2\rho_{1}n]}\nmid b$, and $\vec{v}$ is a primitive vector in the span of $(0,1,0)$ and $(0,0,1)$.
Next, we choose an element $T_{2}\in\text{SL}_{3}(\mathbb{Z})$ such that $T_{2}(X_{0})=(1,0,0)$ and $T_{2}(X_{0})=(0,b,0)$. Let $L_{2}:\mathbb{Z}^{3}\mapsto\mathbb{Z}$ be the linear form getting the third component. Then $L_{2}\circ T_{2}$ is a primitive linear form on $\mathbb{Z}^{3}$. Applying Proposition 4.2 again, for any $Q>q_{4}^{\rho_{1}}$, we have
$\nu^{(l)}(\{g\in\text{SL}_{2}(\mathbb{Z})|L_{2}\circ T_{2}(g\tilde{X}g^{-1})\equiv 0(\operatorname{mod}{Q})\})<Q^{-c_{2}},$
which then implies there is $q_{5}||q_{4}$, $q_{5}>(q_{4})^{\frac{1}{2}}$, and $g_{2}\in A^{l}$ such that for any $p^{n}||q_{6}$, $L_{2}\circ T_{2}(g_{2}X_{0}g_{2}^{-1})\neq 0(p^{[2\rho_{1}n]})$, if
${\delta}<\frac{c_{1}c_{2}\rho_{1}}{96C_{1}}.$ (5.54)
Denoting $V=\text{Lie}(\text{SL}_{2})(\mathbb{Z})$. Multiplying LHS of (5.52) with its $g_{1}$ and $g_{2}$ conjugates, we obtain
Lemma 5.55.
There are $q_{5}||q$, $q_{5}>q^{\frac{c_{1}}{32C_{1}}}$, $q_{5}^{\prime}|q_{5}^{\prime\prime}|q_{5}$, and some constant $K_{2}>0$ depending only on $c_{1}$, such that
$1+q_{5}^{\prime}\mathbb{Z}V\subset A_{1}^{K_{2}}(\operatorname{mod}q_{5}^{\prime\prime}),$ (5.56)
where for each $p^{n}||q_{5}$, $v_{p}(q_{5}^{\prime\prime})=O_{c_{1}}(\rho_{2}n)$ and
$\frac{c_{1}c_{3}}{384C_{1}}\rho_{2}n<v_{p}(q_{5}^{\prime\prime}/q_{5})<v_{p}(q_{5}^{\prime\prime})/2$
if
$\rho_{1}<\frac{c_{1}c_{3}}{3\cdot 2^{9}C_{1}}\rho_{2}.$ (5.57)
5.6. Proof of Proposition 5.1
Collecting all the restrictions (5.5), (5.10), (5.19), (5.29), (5.30), (5.39), (5.50), (5.53), (5.54), (5.57) on $\rho_{1},\rho_{2},\delta$, we find we can take
$\rho_{1}=\frac{c_{1}^{2}\cdot c_{3}}{2^{8}3^{4}C_{1}}\rho_{2}$ (5.58)
$\delta=\frac{c_{1}c_{2}c_{3}}{96C_{1}}\rho_{1}$ (5.59)
We need the elementary lemma whose proof we leave as an exercise:
Lemma 5.60.
Let $H_{1},H_{2}\subset\text{SL}_{2}(\mathbb{Z})$ and let $p$ be a prime. Suppose $1+p^{m_{1}}V(\operatorname{mod}p^{m_{2}})\subset H_{1}(\operatorname{mod}p^{m_{2}})$ and $1+p^{n_{2}}V(\operatorname{mod}p^{n_{2}})\subset H_{2}(\operatorname{mod}p^{n_{2}})$ with $1\leq m_{1}\leq m_{2}\leq 2m_{1}$ and $1\leq n_{1}\leq n_{2}\leq 2n_{1}$. Then
$1+p^{m_{1}+n_{1}}V\subset(H_{1}H_{2})^{4}(\operatorname{mod}p^{m_{2}+n_{2}}).$
Proposition 5.1 is thus proved by iteratively applying Lemma 5.60 to (5.56), with $q_{*}=q_{5}$, $c=\frac{c_{1}}{32C_{1}}$ and $\rho=O_{c_{1}}(\rho_{2})$, $C=O_{S}(\log\frac{1}{\delta})$.
6. A gluing technique for the $\text{SL}_{2}(\mathbb{Z})\times\text{SL}_{2}(\mathbb{Z})$ case
In this section and the next, we prove Proposition 2.2 for the case $\Lambda=\text{SL}_{2}(\mathbb{Z})\times\text{SL}_{2}(\mathbb{Z})$. Suppose all the assumptions about $A$ hold but the conclusion (2.4) does not hold, then by taking $\delta$ sufficiently small, we will find a constant $K$ independent of $\delta$, such that $|\pi_{q}(A^{K})|>|\Lambda_{q}|^{1-\frac{\epsilon}{2}}$. But this will contradict Helfgott’s inequality (5.2) when $\delta$ is sufficiently small.
Proposition 5.1 so far only covers a large subset in one simple factor of the group and for some not too small modulus $q_{*}$. Apply Proposition 5.1 again with $q$ replaced by $q/q_{*}$, we will cover a large subset for some modulus $q_{*}^{\prime}$ coprime to $q_{*}$. This leads to considering the following question: Given a set $B\subset\Gamma_{q_{1}q_{2}},\text{gcd}(q_{1},q_{2})=1$, such that $\pi_{q_{1}}(B),\pi_{q_{2}}(B)$ are large in $\Gamma_{q_{1}},\Gamma_{q_{2}}$, is a bounded power of $B$ covers a large subset of $\Gamma_{q_{1}q_{2}}$?
The following proposition is the main gluing tool.
Proposition 6.1.
Fix $0<\theta<10^{-12}$ and $10<\eta<$ 1, and suppose
$\delta<\frac{c_{2}\eta\theta^{1/2}}{2},$ (6.2)
where $\delta$ is given as in Proposition 2.2 and $c_{2}$ is the implied $c$ in Proposition 4.2. Take
$L>\frac{12}{\delta}.$ (6.3)
Let $q=\prod_{i\in I}p_{i}^{n_{i}}=q_{s}q_{l}$, where $q_{s}=\prod_{p_{i}|q,v_{p_{i}}\leq L}p_{i}^{n_{i}}$ and $q_{l}=\prod_{p_{i}|q,v_{p_{i}}>L}p_{i}^{n_{i}}$. Let $q_{1},q_{2}||q,q_{3}||q_{l}$, $\text{gcd}(q_{1},q_{3})=1$, and $q_{1}q_{2},q_{3}>q^{\eta}$. Suppose for some set $B\subset\Lambda$, we have
$|\pi_{q_{1},q_{2}}(B)|>(q_{1}q_{2})^{3-\theta},$ (6.4)
$|\pi_{q_{3},1}(B)|>q_{3}^{3-\theta}.$ (6.5)
Then there exists $q_{3}^{*}||q_{3}$, $q_{3}^{*}>q_{3}^{\frac{1}{4}10^{-4}}$, such that
$|\pi_{q_{1}q_{3}^{*},q_{2}}(B\cup A)^{O(\log 1/\theta)}|>(q_{1}q_{2}q_{3}^{*})^{3-O(\theta^{\frac{1}{4}})}.$ (6.6)
Remark 6.7.
The point of Proposition 6.1 is we have an increase of modulus from $q_{1}$ to $q_{1}q_{3}^{*}$. To see the necessity of getting help from $A$, for instance, let $q_{1}=1$, $q_{2}=q_{3}$ and $B=\{(\lambda,\lambda),\lambda\in\Gamma\}$. Then $B\cdot B=B$ so there is no hope to expand only from $B$.
Write
$\Gamma_{q_{1}q_{3}}\times\Gamma_{q_{2}}\cong(\Gamma_{q_{1}}\times\Gamma_{q_{2}})\times(\Gamma_{q_{3}}\times 1).$
Since $B$ satisfies (6.4) and (6.5), by Proposition 3.12, there exists $q_{1}^{\prime}|q_{1},q_{2}^{\prime}|q_{2},q_{3}^{\prime}|q_{3},q_{1}^{\prime}q_{2}^{\prime}<(q_{1}q_{2})^{40\theta},q_{3}^{\prime}<(q_{3})^{40\theta}$, such that
$\Gamma(q_{1}^{\prime})/\Gamma(q_{1})\times\Gamma(q_{2}^{\prime})/\Gamma(q_{2})\subset\pi_{q_{1},q_{2}}(B^{2880}).$ (6.8)
$\Gamma(q_{3}^{\prime})/\Gamma(q_{3})\subset\pi_{q_{3},1}(B^{2880}).$ (6.9)
Write
$G=\Gamma(q_{1}^{\prime})/\Gamma(q_{1})\times\Gamma(q_{2}^{\prime})/\Gamma(q_{2})$
From (6.8), we can construct a section map
$\psi:G\rightarrow B^{2880},$
such that $\pi_{q_{1},q_{2}}(\psi(x))=x$.
Since $q_{3}^{\prime}<q_{3}^{40\theta}$, there is $q_{4}||q_{3},q_{4}>q_{3}^{1-\theta^{\frac{1}{2}}}$ such that for every $p^{n}||q_{4}$, we have $p^{[40\theta^{\frac{1}{2}}n]}\nmid q_{3}^{\prime}$.
Write $q_{4}=\prod_{j\in J}p_{j}^{n_{j}}$. For each $p_{j}^{n_{j}}||q_{4}$, we consider $\psi_{j}=\pi_{p_{j}^{[n_{j}\theta^{1/4}]}}\circ\mathbb{P}_{1}\circ\psi$.
Define
$\mathcal{G}_{j}=\{(x,y)\in G\times G|\psi_{j}(xy)\neq\psi_{j}(x)\psi_{j}(y)\}$
According to Proposition 3.1, there are two scenarios:
  1. (1)
    $|\mathcal{G}_{j}|>10^{-4}|G|^{2}.$ (6.10)
  2. (2)
    There is a subset $S_{j}\in G$, $|S_{j}|\geq\frac{99}{100}|G|$ such that $\psi_{j}=h_{j}$ over $S_{j}$ where $h_{j}$ is a homomorphism from $G$ to $\Gamma/\Gamma(p_{j}^{[n_{j}\theta^{\frac{1}{4}}]})$.
Let $J=J_{1}\sqcup J_{2}$ where $J_{1},J_{2}$ is the collection of indices falling into Case (1) and Case (2), respectively. Write $q_{4}=q_{4}^{\prime}q_{4}^{\prime\prime}$, where
$q_{4}^{\prime}=\prod_{j\in J_{1}}p_{j}^{n_{j}},\hskip 28.45274ptq_{4}^{\prime\prime}=\prod_{j\in J_{2}}p_{j}^{n_{j}}.$
We further write $q_{4}^{\prime\prime}=q_{5}q_{5}^{\prime}$, where
$q_{5}=\prod_{\begin{subarray}{c}p_{j}^{n_{j}}||q_{4}^{\prime\prime}\\ \pi_{p_{j}^{[\frac{\theta^{1/4}n_{j}}{2}]}}\circ h_{j}=1\end{subarray}}p_{j}^{n_{j}},\hskip 28.45274ptq_{5}^{\prime}=\prod_{\begin{subarray}{c}p_{j}^{n_{j}}||q_{4}^{\prime\prime}\\ \pi_{p_{j}^{[\frac{\theta^{1/4}n_{j}}{2}]}}\circ h_{j}\neq 1\end{subarray}}p_{j}^{n_{j}}.$
We divide our analysis into three cases.
6.1. The case $q_{4}^{\prime}>q_{4}^{\frac{1}{2}}$
In this case we have
$\sum_{j\in J_{1}}(\log p_{j}^{n_{j}})|\mathcal{G}_{j}|>\log\left((q_{4}^{\prime})^{10^{-4}}\right)|G|^{2}$ (6.11)
The left side of (6.11) is equal to
$\sum_{\begin{subarray}{c}U\subset J_{1}\\ U\neq\emptyset\end{subarray}}\log(\prod_{j\in U}p_{j}^{n_{j}})|\cap_{j\in U}\mathcal{G}_{j}\bigcap\cap_{j\in J_{1}-U}\mathcal{G}_{j}^{c}|$ (6.12)
Since the number of subsets of $J_{1}$ is $<q_{4}^{0+}$, there exists $J_{1}^{\prime}\subset J_{1}$, $\tilde{q}=\prod_{i\in J_{1}^{\prime}}p_{i}^{n_{i}}>(q_{4}^{\prime})^{\frac{1}{2}10^{-4}}$, such that
$|\cap_{i\in J_{1}}\mathcal{G}_{i}|>q^{0-}|G|^{2}.$
Take any $(g_{1},g_{2})\in\cap_{i\in J_{1}^{\prime}}\mathcal{G}_{i}$, and consider $\gamma_{0}=\psi(g_{1})\psi(g_{2})\psi(g_{1}g_{2})^{-1}$. Then $\gamma_{0}$ satisfies,
$\pi_{q_{1},q_{2}}(\gamma_{0})=1,$
and for any $p^{n}||\tilde{q}$,
$\pi_{p^{[n\theta^{{1}/{4}}]}}\circ\mathbb{P}_{1}(\gamma_{0})\neq 1.$
Take a set $B^{\prime}\subset B^{2880}$ which maps bijectively to $\Gamma(q_{3}^{\prime})/\Gamma(q_{3})$ under the projection $\pi_{q_{3},1}$. Take
$B^{\prime\prime}=\{b_{1}\gamma_{0}b_{1}^{-1}\gamma_{0}^{-1}\cdot(b_{2}\gamma_{0}b_{2}^{-1}\gamma_{0}^{-1})\gamma_{0}(b_{2}\gamma_{0}b_{2}^{-1}\gamma_{0}^{-1})^{-1}\gamma_{0}^{-1},b_{1},b_{2}\in B^{\prime}\}$
It follows from a quantitative open mapping result due to Salehi-Golsefidy (see Propostions 49, Proposition 49’ of [Gol19]) that
$\pi_{q_{3}}\circ\mathbb{P}_{1}\left(B^{\prime\prime}\right)\supset\Gamma(q_{3}^{\{O(\theta^{1/4})\}}).$ (6.13)
by first taking the logarithmic map to work on the Lie algebra level, then taking the exponential map back. The definition for $B^{\prime\prime}$ is motivated by the fact that for any $v\in\mathfrak{sl}_{2}(\mathbb{Q})$, $\mathfrak{sl}_{2}(\mathbb{Q})=[\mathfrak{sl}_{2}(\mathbb{Q}),v]+[[\mathfrak{sl}_{2}(\mathbb{Q}),v],v]$, where $+$ here means set addition.
Obviously, $\pi_{q_{1},q_{2}}(B^{\prime\prime})=1$. Multiplying $B^{\prime\prime}$ with $B^{2880}$ and observing (6.8) we obtain Proposition 6.1 in this case by setting $q_{3}^{*}=\tilde{q}$.
6.2. The case $q_{4}^{\prime\prime}>q_{4}^{\frac{1}{2}},q_{5}>(q_{4}^{\prime\prime})^{\frac{1}{2}}$.
The local homomorphisms $h_{j},j\in J_{2}$ can be lifted to a homomorphism
$h:G\rightarrow\Gamma(\prod_{j\in J_{2}}p_{j})/\Gamma((q_{4}^{\prime\prime})^{\{\theta^{\frac{1}{4}}\}}).$
Following the previous reasoning for obtaining $J_{1}^{\prime}$, there is a set $J_{2}^{\prime}\subset J_{2}$, such that
$\bar{q}=\prod_{i\in J_{2}^{\prime}}p_{i}^{n_{i}}>(q_{4}^{\prime\prime})^{\frac{99}{200}},$
$|\cap_{i\in J_{2}}S_{i}|>(q_{1}q_{2})^{-\theta}|G|,$
and $\psi\equiv h$ on $S=\cap_{i\in J_{2}}S_{i}$.
By Proposition 3.12, we have $S^{2880}\supset G^{\prime}=\Gamma(q_{1}^{\prime\prime})/\Gamma(q_{1})\times\Gamma(q_{2}^{\prime\prime})/\Gamma(q_{2})$ for some $q_{1}^{\prime\prime}||q_{1}$, $q_{2}^{\prime\prime}||q_{2}$, $q_{1}^{\prime\prime}q_{2}^{\prime\prime}<(q_{1}q_{2})^{80\theta}$. This implies the existence of a subgroup $G^{\prime\prime}$ of $G^{\prime}$ of the form
$G^{\prime\prime}=\Gamma(\tilde{q}_{1}(q_{1}^{*})^{\{\theta^{\frac{1}{2}}\}})/\Gamma(q_{1})\times\Gamma(\tilde{q}_{2}(q_{2}^{*})^{\{\theta^{{1}/{2}}\}})/\Gamma(q_{2}),$ (6.14)
where
$q_{1}=\tilde{q_{1}}q_{1}^{*},q_{2}=\tilde{q_{2}}q_{2}^{*},\tilde{q}_{1},q_{1}^{*}||q_{1},\tilde{q}_{2},q_{2}^{*}||q_{2},\tilde{q}_{1}\tilde{q}_{2}<(q_{1}q_{2})^{80\theta^{{1}/{2}}}.$ (6.15)
Then we can define a map $\tilde{\psi}:G^{\prime\prime}\rightarrow B^{2880^{2}}$ as, for any $x\in G^{\prime\prime}$, choose a word $x=s_{1}s_{2}\cdots s_{2880}$ and let
$\tilde{\psi}(x)=\psi(s_{1})\psi(s_{2})\cdots\psi(s_{2880}).$ (6.16)
Then we have $\pi_{(q_{4}^{\prime\prime})^{\{\theta^{{1}/{4}}\}}}\mathbb{P}_{1}\tilde{\psi}(x)=h(x)$, and thus $\pi_{(q_{5})^{\{\theta^{{1}/{4}}\}}}\mathbb{P}_{1}\tilde{\psi}(x)=1$.
Recall Lemma 5.43. By iteratively taking commutator of $\tilde{\psi}(G^{\prime\prime})$ $O(\log(1/\theta))$ times, we recover a group
$\Gamma((q_{1}^{*})^{\{{2\theta^{{1}/{4}}}\}}q_{5})/\Gamma(q_{1}^{*}q_{5})\times\Gamma((q_{2}^{*})^{\{{2\theta^{{1}/{4}}}\}})/\Gamma(q_{2}^{*})\subset\pi_{q_{1}^{*}q_{5},q_{2}^{*}}(B^{O(\log(1/\theta))}),$ (6.17)
which is identity when projected to $q_{5}$ and is large when projected to $(q_{1},q_{2}^{\prime})$.
Recall $q_{5}||q_{3}$ and $v_{p}(q_{3}^{\prime})<v_{p}(q_{3})$ for every $p^{n}||q_{5}$. It follows from (6.15), (6.17), and (6.9) that
$|\pi_{q_{1}q_{5},q_{2}}(B^{O(\log 1/\theta)})|>(q_{1}q_{2}q_{5})^{3-{O(\theta^{{1}/{4}})}}.$ (6.18)
In this case, we take $q_{3}^{*}=q_{5}$ and the proposition is proved.
6.3. The case $q_{4}^{\prime\prime}>q_{4}^{\frac{1}{2}},q_{5}^{\prime}>(q_{4}^{\prime\prime})^{\frac{1}{2}}$.
We begin with a lemma.
Lemma 6.19.
Suppose $f$ is a homomorphism from $\Gamma(p_{1}^{m_{1}})/\Gamma(p_{1}^{n_{1}})$ to $\Gamma(p_{2}^{m_{2}})/\Gamma(p_{2}^{n_{2}})$ for some $0\leq m_{1}\leq n_{1},1\leq m_{2}\leq n_{2}$, and for some $\xi\in\Gamma(p_{1}^{m_{1}})/\Gamma(p_{1}^{n_{1}})$, $f(\xi)\neq 1$. Then $p_{1}=p_{2}$ and $p^{x}||\xi-1$ for some $x\leq m_{1}+n_{2}-m_{2}$.
Proof.
We first observe that given $0\leq m_{1}<n_{1},1\leq m_{2}<n_{2}$ and two primes $p_{1},p_{2}$, if there is a nontrivial homomorphism $f$ from $\Gamma(p_{1}^{m_{1}})/\Gamma(p_{1}^{n_{1}})$ to $\Gamma(p_{2}^{m_{2}})/\Gamma(p_{1}^{n_{2}})$, this will force $p_{1}=p_{2}$. To see this, if $m_{1}\neq 0$, then all elements in $\Gamma(p_{1}^{m_{1}})/\Gamma(p_{1}^{n_{1}})$ have orders powers of $p_{1}$, and all elements in $\Gamma(p_{2}^{m_{2}})/\Gamma(p_{2}^{n_{2}})$ have orders powers of $p_{2}$. Since $f$ is nontrivial, we have $p_{1}=p_{2}$.
If $m_{1}=0$, and let us suppose $p_{1}\neq p_{2}$, then $f$ will factor through $\Gamma(p_{1})/\Gamma(p_{1}^{n_{1}})$ and we obtaine a homomorphism $f^{\prime}:\Gamma_{p_{1}}\rightarrow\Gamma(p_{2}^{m_{2}})/\Gamma(p_{2}^{n_{2}})$. Since $\Gamma_{p_{1}}$ is almost simple, $f^{\prime}$ must either be trivial, or injective, or factor thorough the center $Z$. $f^{\prime}$ can not be injective because there are elements in $\Gamma_{p_{1}}$ of order $p_{1}$, which is coprime to $p_{2}$. Likewise, $Z$ can not be the kernel of $f^{\prime}$ because one can also find an element $x\in\Gamma_{p_{1}}/Z$ such that the order of $x$ is coprime to $p_{2}$. Hence, $f^{\prime}$ as well as $f$ must be trivial and we arrive at a contradiction, so $p_{1}=p_{2}$.
Suppose $f(\xi)=y\not\equiv 1(\operatorname{mod}p^{n_{2}})$. Since the order of all elements in $\Gamma(p^{m_{2}})/\Gamma(p^{n_{2}})$ is bounded by $p^{n_{2}-m_{2}}$, it follows that $\xi$ can not be a $p^{n_{2}-m_{2}}$-power in $\Gamma(p^{m_{1}})/\Gamma(p^{n_{1}})$, and this will force $x\leq m_{1}+n_{2}-m_{2}$. ∎
Lemma 6.19 implies that $q_{5}^{\prime}||q_{2}^{*}$, since $(q_{5}^{\prime},q_{1}^{*})=1$.
Let $G^{\prime\prime}$ and $\widehat{\psi}$ be the same as in Case 6.2.
The following isomorphism is clear
$G^{\prime\prime}\cong\Big{(}\Gamma((q_{1}^{*})^{\{\theta^{1/2}\}})/\Gamma(q_{1}^{*})\Big{)}\times\Big{(}\Gamma((q_{2}^{*}/q_{5}^{\prime})^{\{\theta^{1/2}\}})/\Gamma(q_{2}^{*}/q_{5}^{\prime})\Big{)}\times\Big{(}\Gamma((q_{5}^{\prime})^{\{\theta^{1/2}\}})/\Gamma(q_{5}^{\prime})\Big{)}$
Since $\pi_{(q_{5}^{\prime})^{\left\{\theta^{1/4}\right\}}}\circ\mathbb{P}_{1}\circ\tilde{\psi}$ is a homomorphism and factors through $\Gamma((q_{1}^{*})^{\{\theta^{1/2}\}})/\Gamma(q_{1}^{*})\times\Gamma((q_{2}^{*}/q_{5}^{\prime})^{\{\theta^{1/2}\}})/\Gamma(q_{2}^{*}/q_{5}^{\prime})$, it induces a homomorphism
$\widehat{\psi}:\Gamma((q_{5}^{\prime})^{\{\theta^{1/2}\}})/\Gamma(q_{5}^{\prime})\rightarrow\Gamma/\Gamma((q_{5}^{\prime})^{\left\{\theta^{1/4}\right\}}).$
Since $\widehat{\psi}$ is nontrivial after reduction by each exact divisor of $(q_{5}^{\prime})^{\left\{\frac{1}{2}\theta^{1/4}\right\}}$, we can find $g_{1}\in\Gamma((q_{5}^{\prime})^{\left\{\theta^{1/4}\right\}})/\Gamma(q_{5}^{\prime})$ such that for each $p|q_{5}^{\prime}$,
$v_{p}(g_{1})=O(n\theta^{1/4}),\hskip 14.22636ptv_{p}(\widehat{\psi}(g_{1}))<(\frac{1}{2}n\theta^{1/4}).$
Let $g_{2}=g_{1}^{p^{\{\frac{1}{4}n\theta^{1/4}\}}}$, then
$v_{p}(g_{2})=\Theta(n\theta^{1/4}),\hskip 14.22636pt\frac{1}{4}n\theta^{1/4}<v_{p}(\widehat{\psi}(g_{2}))<\frac{3}{4}n\theta^{1/4}.$
Write $g_{2}\equiv 1+Q_{1}X_{1}(\operatorname{mod}Q_{1}^{2}),\widehat{\psi}(g_{2})\equiv 1+Q_{2}X_{2}(\operatorname{mod}Q_{2}^{2})$ for some primitive $X_{1},X_{2}\in\text{Lie}(\text{SL}_{2}(\mathbb{Z}))$. Collecting all powers of $g_{2}$, pulling them back to $G^{\prime\prime}$ and using $\widehat{\psi}$ is induced from $\pi_{(q_{5}^{\prime})^{\theta^{1/4}}}\circ\mathbb{P}_{1}\circ\psi$ by factoring through $\Gamma((q_{1}^{*})^{\{\theta^{1/2}\}})/\Gamma(q_{1}^{*})\times\Gamma((q_{2}^{*}/q_{5}^{\prime})^{\{\theta^{1/2}\}})/\Gamma(q_{2}^{*}/q_{5}^{\prime})$, we obtain
$\pi_{q_{1}^{*}Q_{2}^{\{{4}/{3}\}},(q_{2}^{*}/q_{5}^{\prime})Q_{1}^{2}}(B^{2880^{2}})\cong\left(\pi_{q_{1}^{*},q_{2}^{*}/q_{5}^{\prime}}\times\pi_{Q_{2}^{\{4/3\}},Q_{1}^{2}}\right)(B^{2880^{2}})$
$\supset$ $\Gamma((q_{1}^{*})^{\{\theta^{1/2}\}})/\Gamma(q_{1}^{*})\times\Gamma(({q_{2}^{*}}/{q_{5}^{\prime}})^{\{\theta^{1/2}\}})/\Gamma({q_{2}^{*}}/{q_{5}^{\prime}})\times\left(\pi_{Q_{2}^{\{{4}/{3}\}},Q_{1}^{2}}\left((1,1)+\mathbb{Z}(Q_{2}X_{2},Q_{1}X_{1})\right)\right)$ (6.20)
with $v_{p}(Q_{1}),v_{p}(Q_{2})=\Theta(n\theta^{1/4})$.
With (6.3) at hand, which plays the same role as Theorem B creating a one-parameter group, we run the same arguments in Section 5.5, and we give a sketch here. We first apply Proposition 4.2 with $Q=q_{5}^{\prime}$ to find elements from $A$ to conjugate $X=(X_{1},X_{2})$ to other directions. By considering five primitive linear forms, we can produce $g_{1},g_{2},g_{3},g_{4},g_{5}\in A$ such that the following holds: There exists $\bar{q_{5}}||q_{5}^{\prime}$ and
$\bar{q_{5}}>(q_{5}^{\prime})^{1-5\theta^{\frac{1}{2}}},$ (6.21)
such that
$(\bar{q_{5}})^{\{5\theta^{\frac{1}{2}}\}}\pi_{\bar{q_{5}}}(V\times V)\subset\pi_{\bar{q_{5}}}\text{Span}_{\mathbb{Z}}\{X,g_{1}Xg_{1}^{-1},g_{2}Xg_{2}^{-1},g_{3}Xg_{3}^{-1},g_{4}Xg_{4}^{-1},g_{5}Xg_{5}^{-1}\}$ (6.22)
as long as $\delta<c_{2}\eta\theta^{1/2}/2$.
Multiplying the RHS of (6.3) with its conjugates by $g_{1},g_{2},g_{3},g_{4},g_{5}$, followed by taking commutators of this product iteratively, and finally multiplying all these commutators together, we obtain
$\Gamma((q_{1}^{*}\bar{q_{5}}))^{\{5\theta^{1/2}\}}/\Gamma(q_{1}^{*}\bar{q_{5}})\times\Gamma(((q_{2}^{*}/q_{5}^{\prime})\bar{q_{5}})^{\{5\theta^{1/2}\}})/\Gamma((q_{2}^{*}/q_{5}^{\prime})\bar{q_{5}})$
$\subset$ $\pi_{q_{1}^{*}\bar{q_{5}},(q_{2}^{*}/q_{5}^{\prime})\bar{q_{5}}}\{B\cup A\}^{O(\log 1/\theta)}$
and thus
$|\pi_{q_{1}{q_{5}^{\prime}},q_{2}}\{B\cup A\}^{O(\log 1/\theta)}|>(q_{1}^{*}q_{2}^{*}\bar{q_{5}})^{3(1-5\theta^{1/2})-}>(q_{1}q_{2}q_{5}^{\prime})^{3-O(\theta^{1/2})}$
In this case, we let $q_{3}^{*}={q_{5}^{\prime}}$ and the proposition is proved.
7. Proof of Proposition 2.2 for $\Lambda=\text{SL}_{2}(\mathbb{Z})\times\text{SL}_{2}(\mathbb{Z})$
Let $q=q_{s}q_{l}$, where $q=\prod_{i\in I}p_{i}^{n_{i}}$, $q_{s}=\prod_{i\in I:n_{i}\leq L}p_{i}^{n_{i}}$, $q_{l}=\prod_{i\in I:n_{i}>L}p_{i}^{n_{i}}$ for some $L$ to be determined at (7.12). Let $A$ be given in Proposition 2.2 but we assume
$|\pi_{q}(A\cdot A\cdot A)|<|A|^{1+\delta}.$ (7.1)
We will arrive at a contradiction when $\delta$ is sufficiently small.
We divide our proof into two cases.
7.1. The case $q_{s}>q^{\frac{\epsilon}{2}}$
Take
$\delta<\frac{\log\frac{1}{c_{0}}}{2}$ (7.2)
for $c_{0}=c_{0}(L)$ the implied constant from Theorem A, so that
$|\pi_{q_{s}}^{*}(\chi_{S}^{(l)})(x)-\frac{1}{|\Gamma_{q_{s}}|}|<\frac{1}{2}$
for any $x\in\Gamma_{q_{s}}$. Since $\pi_{q_{s}}^{*}(\chi_{S}^{(l)})(A)>q^{-\delta}$, we have
$|\pi_{q_{s}}(A)|>|\Gamma_{q_{s}}|q^{-\delta}>q_{s}^{6-\frac{2\delta}{\epsilon}-}>q_{s}^{6-2\rho}$ (7.3)
if we take
$\delta<\min\{\frac{\log\frac{1}{c_{0}}}{2},{\epsilon\rho}\}.$ (7.4)
for some $\rho\ll\epsilon$ to be determined at (7.9).
Furthermore, if $q_{s}>q^{1-\frac{\epsilon}{2}}$, then (7.3) implies the assumption in (2.3) is void and so Proposition 2.2 automatically holds. So now we assume $q_{s}<q^{1-\epsilon/2}$.
Now we apply Proposition 5.1 with $q$ replaced by $q_{l}$, then we obtain some $q^{\prime}||q_{l},q^{\prime}>q_{l}^{c}>q^{\frac{\epsilon c}{2}}$ for some $c$ only depending on the generating set $S$, and $C^{\prime}=O_{S}(\log\frac{1}{\rho})$ such that
$|\pi_{q^{\prime}}\circ\mathbb{P}_{1}(A^{C^{\prime}})|>(q^{\prime})^{3-\rho}$
if
$\delta<c^{\prime}\epsilon\rho$ (7.5)
for some $c^{\prime}$ depending only on $S$.
Now we apply Proposition 6.1 with $B=A^{C^{\prime}}$, $q_{1}=q_{2}=q_{s}$, $q_{3}=q^{\prime}$ and $\theta=\rho$. We then obtain $q^{\prime\prime}|q^{\prime},q^{\prime\prime}>(q^{\prime})^{\frac{1}{4}\times 10^{-4}}$, and
$|\pi_{q_{s}q^{\prime\prime},q_{s}}(A^{C^{\prime}C^{\prime\prime}})|>(q_{s}^{2}q^{\prime\prime})^{3-O(\theta^{\frac{1}{4}})},$ (7.6)
for some
$C^{\prime\prime}=O(\log{1}/{\rho}).$ (7.7)
By (7.7), we have increased the modulus of the first component from $q_{s}$ to $q_{s}q^{\prime\prime}$, at the cost of a density loss from $3-\rho$ to $3-O(\rho^{{1}/{4}})$. If $q_{s}q^{\prime\prime}<q^{1-\frac{\epsilon}{2}}$, we apply Proposition 5.1 again with $q$ replaced by $\frac{q}{q_{s}q^{\prime\prime}}$ and $\theta=O(\rho^{1/4})$, after that we can apply Proposition 6.1 again to increase the modulus to a larger one. We apply Proposition 5.1 and Proposition 6.1 iteratively until we reach a modulus $q_{1}^{\star}||q,q_{1}^{\star}>q^{1-\frac{\epsilon}{2}}$. Next, we go through the same procedure to increase the modulus of the second component to $q_{2}^{\star}||q,q_{2}^{\star}>q^{1-\frac{\epsilon}{2}}$. The number of such iterations is bounded by by $\frac{K_{0}\log{(1/\epsilon)}}{c}$ for some absolute constant $K_{0}$.
In the end, we obtain
$\left|\pi_{q_{1}^{\star},q_{2}^{\star}}\left(A^{O_{S}(\log\frac{1}{\rho})}\right)\right|>(q_{1}^{\star}q_{2}^{\star})^{3-O(\rho^{(\frac{1}{4})^{\frac{K_{0}\log(1/\epsilon)}{c}}})}.$ (7.8)
Let $K_{1},K_{2}$ be the implied constants on the LHS and RHS of (7.8), respectively. If we take
$\rho=e^{-(\log(3K_{2}/\epsilon)4^{{\frac{K_{0}\log(1/\epsilon)}{c}}})},$ (7.9)
then from (7.8),
$\left|\pi_{q_{1},q_{2}}\left(A^{K_{1}(\log\frac{1}{\rho})}\right)\right|>q^{6-5\epsilon}.$ (7.10)
Recall also $\pi_{q}(A)<q^{6-6\epsilon}$. Then (7.10) and (5.13) imply
$|\pi_{q}(A\cdot A\cdot A)|>|A|^{1+\frac{\epsilon}{6K_{1}\log(1/\rho)}}.$ (7.11)
We take any $\delta_{0}$ sufficiently small satisfying (7.2), (7.4) and (7.5) for $\delta$. Then we take
$L=\left[\frac{13}{\delta_{0}}\right]$ (7.12)
in view of the condition (6.3). Finally, we set
$\delta=\min\{\delta_{0},\frac{\epsilon}{6K_{1}\log(1/\rho)}\}.$
Then (7.11) contradicts (7.1). Proposition 2.2 is thus proved in this case.
7.2. The case $q_{s}<q^{\frac{\epsilon}{2}}$
In this case the modulus $q_{s}$ can be ignored. We can directly apply Proposition 5.1 and start the iterating process. The analysis is identical to the previous case.
8. Proof of Proposition 2.2 for $\Lambda=\text{SL}_{2}(\mathbb{Z})\ltimes\mathbb{Z}^{2}$
The proof of Proposition 2.2 for the case $\Lambda=\text{SL}_{2}(\mathbb{Z})\ltimes\mathbb{Z}^{2}$ is similar to the case $\Lambda=\text{SL}_{2}(\mathbb{Z})\times\text{SL}_{2}(\mathbb{Z})$. We give a sketch.
We write $q=q_{s}q_{l}$, where $q=\prod_{i\in I}p_{i}^{n_{i}}$, $q_{s}=\prod_{i\in I:n_{i}\leq L}p_{i}^{n_{i}}$, $q_{l}=\prod_{i\in I:n_{i}>L}p_{i}^{n_{i}}$ for some $L$ depending only on the generating set $S$ and $\epsilon$.
We take $\delta$ sufficiently small and assume
$|\pi_{q}(A\cdot A\cdot A)|<|\pi_{q}(A)|^{1+\delta}.$ (8.1)
If $q_{s}>q^{1-\epsilon/2}$, $q_{l}$ can be ignored and we can apply Theorem A to conclude Proposition 2.2.
If $q^{\epsilon/2}<q_{s}<q^{1-\epsilon/2}$, we first apply Theorem A to find some $C_{1}$ so that $A^{C_{1}}(\operatorname{mod}q_{s})$ is large in $\Lambda/\Lambda(q_{s})$. Then we use the method in Section 7 to find some $C_{2}$ so that $\pi_{q,q_{s}}(A^{C_{2}})$ is large in $\text{SL}_{2}(\mathbb{Z}/q\mathbb{Z})\ltimes(\mathbb{Z}/q_{s}\mathbb{Z})^{2}$, and so by Proposition 3.20, there exists $q_{1}||q,q_{2}||(q_{1},q_{s}),q_{1}q_{2}>(qq_{s})^{1-\epsilon/2}$ such that
$\pi_{q_{1},q_{2}}(A^{C_{2}})\supset(\Gamma(q_{1}^{\{\rho\}})/\Gamma(q_{1}))\ltimes(q_{2}^{\{\rho\}}\mathbb{Z}/q_{2}\mathbb{Z})$ (8.2)
for some $\rho=O(\delta)$. Choose $B\subset A^{C_{2}}$ a representative set of RHS of (8.2) and consider the section map
$\psi:(\Gamma(q_{1}^{\{\rho\}})/\Gamma(q_{1}))\ltimes(q_{2}^{\{\rho\}}\mathbb{Z}/q_{2}\mathbb{Z})\rightarrow B.$
If for many $p^{n}||\text{gcd}(q_{l},q_{1})$, $\pi_{p^{\{n\rho^{1/2}\}}}\circ\psi$ is not a homomorphism, then one can apply the argument in Section 6.1 find a large divisor $q_{3}||\text{gcd}(q_{l},q_{1})$ and an element $(g_{0},w_{0})$ from a bounded product of $B$ such that $\pi_{q_{1},q_{2}}((g_{0},w_{0}))=(1,0)$, and $v_{p}(w_{0})=O(\rho n^{1/2})$ for any $p^{n}||q_{3}$. Conjugating $(g_{0},w_{0})$ by $B$ using the formula $(g,x)\cdot(1,y)\cdot(g,x)^{-1}=(1,g(y)),$ and multiplying these conjugates by $B$, we obtain a set which is large in $\Lambda/\Lambda(q_{1})\ltimes(\mathbb{Z}/q_{2}q_{3})^{2}$ under the reduction $\pi_{q_{1},q_{2}q_{3}}$.
If for many $p^{n}||\text{gcd}(q_{l},q_{1})$, $\pi_{p^{\{n\rho^{1/2}\}}}\circ\psi$ is a homomorphism, then we can find a large divisor $q_{4}||\text{gcd}(q_{l},q_{1})$ and a large subset $Z\subset(\Gamma(q_{1}^{\{\rho\}})/\Gamma(q_{1}))\ltimes(q_{2}^{\{\rho\}}\mathbb{Z}/q_{2}\mathbb{Z})$ such that $\psi|_{Z}$ is a homomorphism. But then Proposition 3.24 implies $\psi(Z)(\operatorname{mod}q_{4}^{\{\Theta(\rho^{1/2})\}})\subset H(\operatorname{mod}q_{4}^{\{\Theta(\rho^{1/2})\}})$ for some proper algebraic subgroup $H\subset\text{SL}_{2}(\mathbb{Z})\ltimes\mathbb{Z}^{2}$ defined by two linear equations. Then using Proposition 4.4 and arguing as in Section 5.1 for creating the set $A_{3}$, we can find a large exact divisor $q_{5}||q_{4}$ and $(g_{1},w_{1})$ from a bounded product of $B\cup(A\cdot A)$, such that $\pi_{q_{1},q_{2}}((g_{1},w_{1}))=(1,0)$, and $v_{p}(w_{1})=O(\rho^{1/2}n)$ for any $p^{n}||q_{5}$. With the help of $(g_{1},w_{1})$, we can boundedly generate a set which is large in $\text{SL}_{2}(\mathbb{Z}/q_{1}\mathbb{Z})\ltimes(\mathbb{Z}/q_{2}q_{5}\mathbb{Z})^{2}$.
In either of the above two cases, we increase the modulus $q_{2}$ by a factor of size at least $q^{\kappa}$ where $\kappa>0$ depending only on $S$ and $\epsilon$, Iterating the above procedure by controlled number of times, we can grow $q_{2}$ to a modulus $q_{6}||q_{1}$ and $q_{6}>q_{1}^{1-\epsilon/2}$, which will lead to a contradiction to (8.1) when $\delta$ is sufficiently small.
If $q_{s}<q^{\epsilon/2}$, then $q_{s}$ can be ignored and the expanding procedure is virtually identical to the previous case.